### v1

# MaxMin-RLHF: Towards Equitable Alignment of Large Language Models with Diverse Human Preferences

Abstract
Reinforcement Learning from Human Feedback (RLHF) aligns language models to human preferences by employing a singular reward model derived from preference data. However, such an approach overlooks the rich diversity of human preferences inherent in data collected from multiple users. In this work, we first derive an impossibility result of alignment with single reward RLHF, thereby highlighting its insufficiency in representing diverse human preferences. To provide an equitable solution to the problem, we learn a mixture of preference distributions via an expectation-maximization algorithm and propose a MaxMin alignment objective for policy learning inspired by the Egalitarian principle in social choice theory to better represent diverse human preferences. We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution. We present comprehensive experimental results on small-scale (GPT-2) and large-scale language models (with Tulu2-7B) and show the efficacy of the proposed approach in the presence of diversity among human preferences. Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach. We remark that our findings in this work are not only limited to language models but also extend to reinforcement learning in general.
1 Introduction
The alignment problem, central to developing and fine-tuning current large language models (LLMs), represents a crucial challenge in artificial intelligence, especially in ensuring these models operate in harmony with human values, preferences and social welfare (Wang et al., 2023; Christian, 2020). Reinforcement learning from human feedback (RLHF) has emerged as a pivotal approach to alignment problems, specifically aligning LLM (Wang et al., 2023; Ouyang et al., 2022b; Stiennon et al., 2022a; Ouyang et al., 2022a). RLHF operates in three steps (a) supervised fine-tuning, (2) reward learning, and (3) RL fine-tuning. Step 2 learns a reward function that is expected to represent the preference feedback of the human population. However, there has been minimal emphasis on accurately representing the diversity of human preferences and the broad spectrum of user populations. As highlighted by Aroyo & Welty (2015); Aroyo et al. (2023a, b), “the notion of ‘one truth’ in crowdsourcing responses is a myth” and we need to account for the diversity in opinions and preferences.
Despite the criticality, most of the latest RLHF approaches ignore the consideration of the diversity in human preference feedback by aligning the language model with a single reward function (Wang et al., 2023; Christian, 2020; Stiennon et al., 2022a; Ouyang et al., 2022a). This assumption of a single ground truth reward is restrictive and can potentially subdue the preferences or opinions of minority groups, leading to societal biases (Figure 1). To mitigate this issue, some of the recent research proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners (Bakker et al., 2022; Jang et al., 2023). On the other hand, (Ovadya, 2023) adopts a consensus-based method for aggregating human representations by emphasizing specific principles (Bai et al., 2022b; Kovač et al., 2023), which might result in the under-representation of marginalized groups (Ramé et al., 2023). Another line of research focuses on the aspect of designing multi-policy strategies by fine-tuning personalized language models towards individual rewards (Jang et al., 2023; Ramé et al., 2023; Ji et al., 2023a).
Some recent literature has brought attention to the challenge of aligning single utility RLHF with diverse preferences (Bakker et al., 2022; Ramé et al., 2023; Jang et al., 2023). However, a thorough understanding of how the diversity within human sub-populations influences the overall alignment objective remains elusive. Consequently, this prompts us to pose the following question:
Is the current single reward RLHF pipeline sufficient to align with diverse human preferences?
In this work, we present negative results for the above question by demonstrating the impossibility of alignment using single reward RLHF (Theorem 3.3). We introduce a notion of diversity between human subpopulations due to the differences in preference distributions and establish lower bounds on the alignment performance of single reward RLHF. However, this impossibility result naturally raises another important question:
What strategies can we design to efficiently align with diverse human preferences?
In response to this question, we draw inspiration from the Egalitarian rule (Sen, 2017) to provide an equitable solution to the problem and aim to maximize the social utility objective for alignment. We propose learning a mixture of preference distributions via expectation-maximization algorithm in absence of human sub-population group information. Under the mixture distributions, we introduce the notion of MaxMin RLHF, for policy learning inspired by the Egalitarian principle in social choice theory to better represent human preferences under diversity. We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution. We summarize our contributions as follows.
Figure 1: This figure highlights the drawbacks of a single reward-based current state-of-the-art alignment framework called Reinforcement Learning from Human Feedback (RLHF) (Christian, 2020). In this figure, we demonstrate a setting where, due to the inherent presence of majority and minority user groups who provide human feedback, single reward-based RLHF alignment would align the language model towards the majority group while completely ignoring the minority use group preferences. We provide a theoretical justification in Section 3 and empirical evidence in Section 5.
(1) An impossibility result of alignment with single reward-based RLHF. We first introduce the notation of diversity as a divergence between the human sub-population preference distributions (Definition 2) and derive lower bounds on the reward model sub-optimality (Lemma 3.2) in terms of the diversity. Finally, we establish a lower bound (Theorem 3.3) on the alignment gap due to the diversity in the human preference feedback. True to our knowledge, ours is the first work to formally characterize the sub-optimality gap with conventional RLHF methods under diverse preferences.
(2) Max-Min RLHF alignment with diverse user preferences. To provide an equitable solution to the problem, we learn a mixture of preference distributions using the Expectation-Maximization (EM) algorithm (Algorithm 2) and propose a novel MaxMin-RLHF algorithm as an egalitarian strategy to align language models with social utility objectives under diverse preferences. (Algorithm 1).
(3) A comprehensive empirical study. We present a detailed empirical analysis of our proposed approach on both small (GPT-2) and large language models (Tulu-7B). Initially, we provide empirical evidence highlighting the impossibilities of alignment with single reward RLHF, followed by demonstrating the feasibility and effectiveness of MaxMin-RLHF in achieving social utility objectives. Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach.
1.1 Related Works
The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017). Besides PPO, DPO (Direct Preference Optimization, Rafailov et al. (2023)) directly trains the large language model using human preferences without training the reward model. A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework. However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023). Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a). The key factors contributing to this diversity include (i) socio-demographic backgrounds (Vogels, 2021), (ii) personal bias and context subjectivity (Denton et al., 2021b; Sandri et al., 2023)), (iii) Imperfect preferences, (Sandri et al., 2023), and (iv) Linguistic ambiguity & missing context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022). These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback. Appendix B provides a detailed context of related works.
2 Preliminaries
Let us start by defining a language model mathematically. We denote a vocabulary set as $\mathcal{V}$ and a language model by a mapping $\pi_{\theta}$ (parameterized by $\theta$). A language model $\pi_{\theta}$ takes a sequence of tokens (called prompt) as input denoted by $\mathbf{x}:=\{x_{1},x_{2},\cdots,x_{N}\}$, where each token $x_{i}\in\mathcal{V}$. The prompt $\mathbf{x}\in\mathcal{X}$, where $\mathcal{X}$ is the set of prompts, is fed as input to the language model, and it generates output response $\mathbf{y}\sim\pi_{\theta}(\cdot~{}|~{}\mathbf{x})$.
RLHF pipeline. We start by considering the RLHF pipeline in Ziegler et al. (2019), which has also been adopted in subsequent works (Stiennon et al., 2022c; Bai et al., 2022a; Ouyang et al., 2022b). It consists of three steps detailed as follows:
Step 1: Supervised Fine-tuning (SFT): In this phase, a generic pre-trained LM is fine-tuned with supervised learning on a high-quality dataset for the downstream task(s) of interest, such as dialogue, instruction following, summarization, etc., to obtain a model $\pi_{\text{ref}}$.
Step 2: Reward Modelling: In the second phase, the SFT model is queried with prompts $\mathbf{x}\in\mathcal{X}$ to produce pairs of responses $(\mathbf{y}_{1},\mathbf{y}_{2})\sim\pi_{\theta}(\cdot~{}|~{}\mathbf{x})$ which are then presented to human labelers for preference evaluation, and $\mathbf{y}_{1}$, $\mathbf{y}_{2}$ denotes the preferred and dispreferred response, respectively. The preference distribution under the Bradley-Terry (BT) preference model (Bradley & Terry, 1952) is written as
$p^{*}(\mathbf{y}_{1}\!\succ\!\mathbf{y}_{2}\mid\mathbf{x})\!=\!\frac{\exp\left(r^{*}(\mathbf{y}_{1},\mathbf{x})\right)}{\exp\left(r^{*}(\mathbf{y}_{1},\mathbf{x})\right)+\exp\left(r^{*}(\mathbf{y}_{2},\mathbf{x})\right)},$ (1)
where $r^{*}(\mathbf{y},\mathbf{x})$ is the latent reward model. With a static dataset $\mathcal{D}=\bigl{\{}\mathbf{x}^{(i)},\mathbf{y}_{1}^{(i)},\mathbf{y}_{2}^{(i)}\bigr{\}}_{i=1}^{N}$ sampled from $p^{*}$, we can learn a parameterized reward model $r_{\phi}(\mathbf{y},\mathbf{x})$ via maximum likelihood estimation. Framing the problem as a binary classification, we have the negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})\!=\!-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\bigl{[}\log\sigma(r_{\phi}(\mathbf{y}_{1},\mathbf{x})-r_{\phi}(\mathbf{y}_{2},\mathbf{x}))\bigr{]}$ (2)
where $\sigma$ is the logistic function.
Step 3: RL Fine-Tuning: In the final step, the optimal policy $\pi_{r_{\phi}}^{*}$ under the reward $r_{\phi}$ is obtained by solving the KL-regularized reward maximization problem given by
$\max_{{\pi}}\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot|\mathbf{x})||\pi_{\text{ref}}(\cdot|\mathbf{x})\bigr{]}\bigr{]},$ (3)
where, $\beta>0$ controls the deviation from the base reference policy $\pi_{\text{ref}}$.
3 An Impossibility Result for Single Reward RLHF with Diverse Preferences
In this section, we mathematically prove the impossibility of aligning language models with diverse human preferences with the single reward RLHF framework. We start by discussing the motivation and mathematical definition of diversity in human preferences in Section 3.1, then connect the reward learning step of the RLHF pipeline with diversity in Section 3.2, and then finally prove the impossibility of language model alignment in Section 3.3 by connecting Step 3 of RLHF pipeline with human preference diversity.
3.1 Diversity in Human Preferences
Figure 2: (Diversity in Preferences.) This figure illustrates the diversity in preferences among two distinct human groups using the IMDB movie review dataset (Maas et al., 2011). We categorize these groups as ‘majority’ and ‘minority.’ (a) and (c) display minority sentiment and conciseness preferences. We note that the minority group strongly favors concise responses (as seen in the blue curve in (c)), while showing indifference towards sentiment (as indicated by overlapping curves in (a)). In contrast, (b) and (d) depict that the majority clearly prioritizes positive sentiment (as evidenced by a significant gap between chosen and rejected trajectories in (b)), while displaying little concern for conciseness (as indicated by overlapping curves in (d)).
The main shortcoming of state-of-the-art alignment approaches arises from the underlying assumption that human preferences are derived from a single latent reward model $r^{*}(\textbf{y},\textbf{x})$ (cf. (2)), which fails to account for the inherent diversity among the human sub-populations (see Figure 2). As discussed in Section 1.1, one of the key reasons for the diverse human preferences is the varied socio-demographic and socio-cultural backgrounds of human sub-populations (Aroyo et al., 2023b, a). For example, population groups with diverse demographic markers such as race, ethnicity, age groups, genders, etc., have highly varied preferences as highlighted in (Aroyo et al., 2023b, a; Denton et al., 2021a). Such diversity inevitably leads to natural sub-groups of populations among humans. Modeling this diversity in preferences for the fine-tuning of language models in RLHF is crucial, which, to the best of our knowledge, is currently missing from the literature.
Sub-population Preference Distributions: Let us consider the human population providing the preference feedback represented by $\mathcal{H}$. We can write the preference distribution (Stiennon et al., 2022a; Ouyang et al., 2022a) as
$p^{*}($ $\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ (4)
[…]
where $p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ is the probability of preferring ${\mathbf{y}}_{1}$ over ${\mathbf{y}}_{2}$ for any given pair $(\mathbf{y}_{1},\mathbf{y}_{2})$ corresponding to prompt $\mathbf{x}$. In (4), the expectation is over a finite set of humans $h\in{\mathcal{H}}$. We next introduce the concept of human subpopulations as a hidden random variable, denoted as $u$ with distribution $\eta$, to account for the inherent diversity within the population. Specifically, $u$ represents the human subpopulation defined over a finite discrete set $\mathcal{U}:=\{\mathcal{H}_{1},\mathcal{H}_{2},\cdots,\mathcal{H}_{|\mathcal{U}|}\}$, such that $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$. The cardinality of the set $\mathcal{U}$ represents the number of sub-populations/groups present in the total human population $\mathcal{H}$. Therefore, similar to (4), we can define a human-subpopulation or group-specific preference distribution for a given pair of responses $(\mathbf{y}_{1},\mathbf{y}_{2})$ and prompt $\mathbf{x}$ as
$p^{*}_{u}($ $\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ (5)
[…]
for all groups in $\mathcal{U}$. Next, we define the preference diversity among the human population in Definition 2 as follows.
{defi}
[Diversity in Human Preferences] Consider a human population $\mathcal{H}$, composed of $|\mathcal{U}|$ sub-population groups where $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$, and a sub-population-specific preference $p^{*}_{u}$ as defined in (5), we define the diversity of sub-population group $\mathcal{H}_{i}$ with respect to other group $\mathcal{H}_{j}$ as
$\texttt{Diverity}\ (i,j):=\text{TV}(p_{i}^{*},p^{*}_{j}),$ (6)
where TV denotes the total variation distance between two preference distributions. By utilizing the definition of sub-population groups in $\mathcal{U}$, we can express the preference in (4) as
$p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=$ $\sum_{u=1}^{|\mathcal{U}|}\bigg{[}\sum_{h\in\mathcal{H}_{u}}\mathbb{I}_{h}(\mathbf{z})\cdot q(h|u)\bigg{]}\cdot\eta(u)$
$=$ $\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{z}})\cdot\eta(u),$ (7)
where $\mathbf{z}:=(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ is a shorthand notation and $q(\cdot)$ denotes the distribution over the humans $\mathcal{H}$. Here, $p_{u}^{*}({\mathbf{z}}_{h})=\sum_{h\in\mathcal{H}_{u}}\mathbb{I}_{h}(\mathbf{z})\cdot q(h|u)$ is the sub-population specific preference distribution (cf. (5)) and $\eta(\cdot)$ represents the marginal probability distribution of sub-population $\mathcal{H}_{u}$ and quantifies the probability of occurrence of sub-population $\mathcal{H}_{u}$ to provide feedback for pair z. We can think of $\eta(\cdot)$ as a weighting function that quantifies the relative importance of each sub-population (say $\mathcal{H}_{u}$) within the full population $\mathcal{H}$ reflecting their contributions to the aggregate preference distribution $p^{*}$. Thus, from the expansion in (3.1), it is evident that the preference distribution under consideration is a weighted sum of sub-population specific preference distribution, weighted by $\eta(u)$. We remark that distributions $q$ and $\eta$ are crucial to rigorously characterize the alignment performance of different approaches, which is not considered in the existing literature (Christian, 2020; Bai et al., 2022a).
3.2 Reward Mismatch Due to Diversity
From equations (1) and (2), we note that the existing RLHF approach focuses on learning the ground-truth single reward parameter $\phi^{*}$ to represent the preference distribution $p^{*}$ by minimizing the cross-entropy loss (cf. (2)) given by
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=$ $-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\Big{[}p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})\log p_{\phi}(\succ)$
$+p^{*}(\mathbf{y}_{1}\prec\mathbf{y}_{2}\mid\mathbf{x})\log p_{\phi}(\prec)\Big{]},$ (8)
The assumption of single ground-truth reward (corresponding to $p^{*}$) which is violated due to the existence of diverse sub-populations with separate preference distributions, as discussed in Section 3.1. This would lead to an implicit aggregation as shown in (3.1) and the equivalent MLE objective in (3.2) can be re-written as :
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})$ $\!=\!-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\Big{[}\mathbb{E}_{u}[p^{*}_{u}(\mathbf{y}_{1}\!\succ\!\mathbf{y}_{2}\mid\!\mathbf{x})]\log p_{\phi}(\succ)$
$+\mathbb{E}_{u}[p^{*}_{u}(\mathbf{y}_{1}\prec\mathbf{y}_{2}\mid\mathbf{x})]\log p_{\phi}(\prec)\Big{]}.$ (9)
Now, expanding upon the cross-entropy objective, we note (see Lemma C for details) that the objective in (3.2) essentially reduces to minimizing the Kullback-Leibler (KL) divergence $\mathsf{KL}(\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{u}^{*}({\mathbf{z}})||p_{\phi})$ and the objective is minimized at $p_{\phi^{*}}=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{u}^{*}$. This implies that by minimizing the loss function in (3.2), when we try to learn a single $\phi^{*}$ to recover $p^{*}$, an implicit averaging happens over the preferences of human subpopulation groups they belong to, which plays a critical role in the sub-optimality in reward learning summarized in Lemma 3.2. {lem}[] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf. 1) and $\phi^{*}_{u}$ models the human sub-population group $\mathcal{H}_{u}\in\mathcal{U}$ specific $p_{u}^{*}$, it holds that
$\underbrace{\|\phi^{*}-\phi_{u}^{*}\|}_{\textbf{Reward mismatch}}\geq\frac{1}{2D}\cdot{\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\texttt{Diverity}\ (u,k)},$
where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, and diversity as defined in Definition 2. Proof Sketch. Here we describe the proof sketch of Lemma 3.2 with a detailed proof provided in Appendix D. We begin with the definition of sub-optimality in the learned reward for a subpopulation group $u$ as $\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*}$ where $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter $\phi^{*}$ which minimizes the loss function in (2). However, we know in the limit of infinite data, $\hat{\phi}_{\texttt{MLE}}$ converges to $\phi^{*}$ and hence we focus on the sub-optimality gap due to diversity as $\|\phi_{u}^{*}-\phi^{*}\|$. Using the Lipschitzness of the preference probability distribution under the Bradley-Teryy preference model (derived in Lemma C in Appendix) we lower-bound the sub-optimality gap by $\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ and finally expanding upon the definition of $p^{*}$ as shown in (3.1), we get the final result.
Remark. Lemma 3.2 indicates that the current RLHF-based reward learning paradigm (Christian, 2020; Bai et al., 2022a; Rafailov et al., 2023) will suffer sub-optimality due to diversity amongst the humans, which is highly likely in practice (Aroyo et al., 2023b). Lemma 1 implies that the degree to which the learned reward parameter diverges from optimality for a given subgroup is influenced by two key factors: the distinctiveness of that subgroup’s preferences compared to all the other subgroups, and the relative weight assigned to the subgroup in the overall preference model.
3.3 An Impossibility Result of Alignment
To mathematically characterize the impossibility of aligning the language model with diverse sub-population groups, let us reconsider the RL fine-tuning optimization problem, which is given by (step 3 in RLHF in Section 2)
$\max_{{\pi}}F_{{\phi}}(\pi),$ (10)
Figure 3: (Empirical Evidence of Impossibility). This figure provides empirical evidence of the impossibility of alignment (cf. Theorem 3.3) in single reward RLHF on preference dataset presented in Figure 2. Here, the task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature. The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right). The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right). We note that the aligned language model (top row right) can generate highly positive sentiment sentences but completely ignores the requirement of conciseness (generates longer sentences). This is happening because the humans who prefer conciseness are in minority as described in Figure 2.
where we define $F_{{\phi}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}\Big{[}\mathbb{E}_{\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})\bigr{]}-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot~{}|~{}\mathbf{x})||\pi_{\text{ref}}(\cdot~{}|~{}\mathbf{x})\bigr{]}\Big{]}$. Let us define $\pi_{\text{RLHF}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}}}(\pi)$ where $\pi_{\text{RLHF}}$ is the optimal aligned policy with single reward RLHF. On the other hand, we define a human sub-population specific optimal policy as $\pi^{*}_{\text{u}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}_{u}}}(\pi)$, where $\pi^{*}_{\text{u}}$ is the optimal aligned policy with individual subpopulation group $\mathcal{H}_{u}$. We define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $\mathcal{H}_{u}$ by
[…]
We note that the alignment gap defined in (11) measures the discrepancy between the reward returns by the single reward RLHF model $\pi_{\text{RLHF}}$ and the optimal model $\pi_{u}^{*}$ tailored for $\mathcal{H}_{u}$ subpopulation evaluated under true reward function $r_{u}^{*}$. Next, we present our impossibility result in Theorem 3.3. {thm}[An Impossibility Result] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf. 1), $\phi^{*}_{u}$ denotes the human sub-population group $\mathcal{H}_{u}\in\mathcal{U}$ specific reward function to model $p_{u}^{*}$, and alignment gap is as defined in (11). Then, it holds that
[…]
where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, $\lambda_{\psi}$ denotes the minimum eigenvalue of the feature matrix, $\beta$ is the regularization parameter of RLHF framework, and diversity as defined in Definition 2. A detailed proof of Theorem 3.3 is provided in Appendix E. We briefly describe the proof sketch of Theorem 3.3 as follows.
[…]
Remark. Theorem 3.3 shows that high subpopulation diversity inevitably leads to a greater alignment gap for the model $\pi_{\text{RLHF}}$ aligned with single reward-based RLHF framework. In summary, if a subgroup exhibits distinctive preferences or constitutes a minority with a smaller representation, the resulting model from single reward RLHF setting cannot accurately reflect the sub-population’s specific preferences. We provide empirical evidence of impossibility of alignment in Figure 6.
4 Proposed Approach: Alignment with Diverse Preferences
From the statement of Theorem 3.3, we note that it is impossible to align the language model with diverse human preferences with a single reward model based RLHF. Now the natural question arises is what can we do in practice to deal with the diverse human preferences alignment problem? This is exactly what we answer in this section. We start by proposing to learn a mixture of preference distributions as follows.
Algorithm 1 MaxMin RLHF
1:  Input: Preference dataset $\mathcal{D}$, initial reward parametrization for each subpopulation $u$ as $r_{\phi_{0}}^{u}$, initial policy parameter $\pi_{0}$.
2:  Reward Learning with EM: Utilize Algorithm 2 for learning rewards with EM to learn $r_{\phi}^{u}$ for all user subpopulation $u$
3:  Max-Min Policy Iteration:
4:  for $t=0$ to $T-1$ do
5:     Choosing Minimum Utility Subpopulation:
6:     $u_{\text{min}}\leftarrow\arg\min_{\mathcal{H}_{u}\in\mathcal{U}}F_{r_{\phi}^{u}}(\pi_{t})$
7:     Perform the PPO Update:
8:     Update policy $\pi$ towards maximizing the objective:
9:     $\pi_{t+1}\leftarrow\text{PPO-update}(F_{{\phi_{u}^{*}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]})$
10:  end for
11:  Output: Policy $\pi_{T}$ aligned with socially fair preference dataset
4.1 Learning a Mixture Distribution for Diverse Preferences
To deal with the diversity in preferences, we first propose learning a mixture distribution for preferences rather than a single preference distribution, which is prevalent in the current RLHF pipelines (Christian, 2020; Ouyang et al., 2022b). The intuition behind learning a mixture of preference models is rooted in the principles of statistical learning theory (Lindsay, 1995; Seidel, 2011) which states that a combination of multiple distributions can approximate a wide range of complex, real-world distributions more effectively than a single model. This concept is akin to the property for mixture models, which suggests that a mixture model can approximate any number of diverse user sub-populations to an arbitrary level of accuracy, given a sufficient number of components and appropriate parameterization (Goodfellow et al., 2016). Thus, the preference distribution as shown in (3.1) can be expressed as
$p({\mathbf{z}}^{\prime})=\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (13)
where the event $\mathbf{{\mathbf{z}}}^{\prime}:=(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})$, $y_{w},y_{l}$ are the chosen and rejected response by the human sub-population-group $\mathcal{H}_{u}$. With the definition in (13), the log-likelihood maximization objective to learn the sub-population preference distributions $\{p_{\phi_{u}}\}_{u=1}^{|\mathcal{U}|}$ parametrized by $\Phi=[\phi_{1},\phi_{2},\cdots\phi_{|\mathcal{U}|}]$ (under the Bradley Terry preference model in (1)) and the $|\mathcal{U}|$ groups/clusters can be written as
$\mathcal{L}(\Phi)$ $=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (14)
$=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}\frac{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)}{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)+\exp\left(r_{\phi_{u}}(\mathbf{y}_{l},\mathbf{x})\right)}\cdot\eta(u).$
We learn the parameter $\Phi^{*}$ by maximizing the above log-likelihood objective. It is important to emphasize that mixture models do not require prior knowledge of user group classifications, i.e., which user belongs to which group, and can learn the distribution of preferences in an unsupervised manner. This capability underscores the generality and flexibility of mixture distributions in capturing the diversity of user preferences without explicit user group information.
Remark. If we have access to individual human sub-population information, we can directly learn the reward parameter via maximizing likelihood per user group. But often, in practice, they are hardly available. To address this challenge, we consider an expectation-maximization algorithm to learn a mixture of reward models summarized in Algorithm 2 which learns the $r_{\phi_{u}}$’s and the $|\mathcal{U}|$ clusters. We summarize the EM algorithm for reward learning in Algorithm 2.
Figure 4: (a) This figure shows a Robotic navigation scenario where the objective of the agent is to distribute goods among the two groups denoted by green and orange boxes. In the above figure, (b) shows the trajectory when only green user group preferences (they want the agent to visit them) are considered to decide the agent path. (c) shows the trajectory when only green user group preferences are considered to decide the vehicle path. (d) shows the result for our proposed formulation (cf. Definition 4.2), where goal is to maximize the social utility, which ensures a robust solution to satisfy all user group preferences.
4.2 MaxMin RLHF
We begin by noting that while we circumvent the sub-optimality in reward learning (cf. Lemma 3.2) sub-population specific reward models ${\phi}^{*}_{u}$ for all subsets in $\mathcal{U}$ using Algorithm 2, this does not address the ultimate goal of language model alignment. Since the eventual objective is to develop a single policy $\pi^{*}$ that accommodates diverse user preferences while avoiding bias towards specific groups such as minorities. To achieve that, we take motivation from the Egalitarian rule in social choice theory (Sen, 2017), which states that society should focus on maximizing the minimum utility of all individuals. Hence, we write our proposed alignment objective, which maximizes the social utility as follows {defi}(Alignment with Diverse Human Preferences) From the mixture of preference distributions in (14), we get $\{r_{\phi_{u}^{*}}\}_{u=1}^{|\mathcal{U}|}$ where $\phi_{u}^{*}$ denotes reward model parameter for each human subpopulation in $\mathcal{U}$. We define the alignment objective in such diverse settings as
$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\Big{(}\min_{u}\mathbb{E}\big{[}{r_{\phi_{u}^{*}}}(\mathbf{y},\mathbf{x})\big{]}\Big{)}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (15)
where the expectation is with respect to $\{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})\}$.
Proof of Concept: An illustration of MaxMin RLHF. To illustrate the social utility of the Max-Min objective within the context of RLHF, in Figure 4, we present a simplified robotic scenario. In this scenario, the robotics agent’s goal is to distribute goods among two user groups $u_{1},u_{2}$ located at green and orange boxes, respectively. Each user group would prefer that the agent visit them and distribute goods. Hence, we define the reward for each user group $u_{i}$ is defined as follows: the agent receives a reward $r_{u}(s,a)=C_{u}$ if it reaches the goal state for the user $u$ and $r_{u}(s,a)=0$, otherwise. We demonstrate that in this basic scenario, our proposed max-min algorithm ensures that the agent serves all the user groups irrespective of the difference in utility between the user groups $\Delta_{C}=|C_{1}-C_{2}|$. Mathematically, from the Definition 4.2, we can write
$\pi^{*}_{\text{MaxMin}}\in\mathop{\mathrm{argmax}}_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (16)
[…]
$\pi^{*}_{\text{MaxMin}}\in$ $\mathop{\mathrm{argmax}}_{\pi}\min_{u}\sum_{s,a}r_{\phi_{u}}(s,a)\cdot\lambda^{\pi}(s,a)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
$\in$ $\mathop{\mathrm{argmax}}_{\pi}{\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (17)
where $\lambda^{\pi}$ denote the occupancy measure corresponding to policy $\pi$. Let us focus on the minimization objective $\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})$, where $(s_{u},a_{u})$ signifies the goal state and action for a user $u$, and $\mathcal{C}_{u}$ is the corresponding reward. When the policy $\pi$ does not visit the goal for user $u$, the value of $\lambda^{\pi}(s_{u},a_{u})$ is zero, making the minimum always zero. Therefore, the maximum objective aims to learn a policy that maximizes the reward with respect to the users $u$ until $\lambda_{u}>0$ for all users. Even when the difference $\Delta_{c}=|\mathcal{C}_{1}-\mathcal{C}_{2}|$ is significantly large, highlighting a disparity between majority and minority user groups, the max-min approach effectively learns a policy that respects the needs of both users, demonstrating its social utility as shown in Figures 4.
MaxMin RLHF Algorithm. Next, we introduce the algorithm to solve the MaxMin RLHF proposed in Definition 4.2 in Algorithm 1. After learning the reward in Step 1 for individual user sub-population $r_{u}$ using Algorithm 2, we next compute the minimum utility user group as $u_{\text{min}}\leftarrow\arg\min_{u}F_{u}(\pi)$. Once the minimum utility user subpopulation group is determined, we fine-tune the language model policy for that group by taking gradient steps at each iteration $t$ as
$\pi_{{t+1}}=\pi_{t}+\eta\nabla_{\pi}(F_{u_{\text{min}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]}),$ (18)
where $F_{u_{\text{min}}}(\pi_{t}):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi_{t}(\cdot~{}|~{}\mathbf{x})}\big{[}{r_{u_{\text{min}}}}(\mathbf{y},\mathbf{x})\big{]}$ and $\eta$ denotes the step size. Note that for the implementation in practice, we leverage the PPO algorithm (Schulman et al., 2017) as commonly used in existing RLHF pipelines (Stiennon et al., 2022b; Christian, 2020; Ouyang et al., 2022b).
Algorithm 2 Learning Mixture Distribution with EM Algorithm
1:  Input: Preference data $\mathcal{D}$, $|\mathcal{U}|$ clusters of users among all humans in $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$, pretrained $\{r_{\phi_{u}}\}_{u=1}^{|\mathcal{U}|}$, loss function $\operatorname{loss}$, convergence criteria
2:  while not reach the convergence criteria do
3:     for $h\in\mathcal{H}$ do
4:        E-step (hard cluster assignment): assign $h$ to the $u$-th cluster s.t.
$\mathcal{H}_{u}:=\ arg\max_{u\in{1,\cdots,|\mathcal{U}|}}\prod_{(\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},h)\in\mathcal{D}}w(\phi_{u},\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}})$
where $w(\cdot)=\frac{\exp\left(r_{\phi_{u}}(\mathbf{y_{1}},\mathbf{x})\right)}{\exp\left(r_{\phi_{u}}(\mathbf{y_{1}},\mathbf{x})\right)+\exp\left(r_{\phi_{u}}(\mathbf{y_{2}},\mathbf{x})\right)}$
5:     end for
6:     M-step: Update each $\phi_{u},u=1,\cdots,|\mathcal{U}|$ by minimizing the negative log-likelihood loss (14) on the assigned users’ data
7:  end while
4.3 Connection with DRO and General Utility RL
In this subsection, we highlight the generality of our proposed formulation beyond social utility design to broader formulations of general utility reinforcement learning (Zhang et al., 2020) and distributionally robust optimization (Duchi et al., 2018). We discuss then in details follows.
1. Distributionally Robust Optimization: We begin by noting that once we have learned the mixture of reward models $r_{{\phi}_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the problem of learning a robust policy as a distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Duchi et al., 2018). In DRO, the objective is to maximize the worst-case expected utility over an uncertainty set of distributions. In our scenario, this boils down to constructing an uncertainty set over the distribution of rewards $r\in\mathcal{Q}$. Now, expanding the RL fine-tuning objective in the equation
$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\min_{r\in\mathcal{Q}}\mathbb{E}_{r(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (19)
where $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3). The uncertainty set $\mathcal{Q}$ is typically obtained by constructing a distribution across the reward distributions $r$, however, this may lead to an overtly pessimistic solution. Hence, recent literature (Sagawa et al., 2020) has shown leveraging the prior knowledge of user-group information, can mitigate this by constructing the uncertainty set $\mathcal{Q^{\prime}}$ in terms of these groups $\mathcal{H}_{u}\in\mathcal{U}$, also referred to as the group distributionally robust optimization (Group-DRO). In Group DRO, the uncertainty set over reward distribution $\mathcal{Q}$ is constructed by a mixture of $u$ groups as $\mathcal{Q^{\prime}}=\sum_{u}\delta_{u}r_{u},\delta_{u}\in\Delta_{u}$, where $\Delta_{u}$ represents the $u-1$ dimensional probability simplex of preference distributions. However, since the inner problem is a linear program and the optimum is attained at the vertex, the resulting group DRO problem boils down to solving
$\pi^{*}_{\text{MaxMin}}$ $\in\arg\max_{\pi}\min_{u}\mathbb{E}_{r_{u}(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (20)
$\in\arg\max_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
where $r_{u}(y,x)$ represents the reward distribution of the user sub-population group $u$. It is important to note that the group-DRO formulation boils down to the Max-Min RLHF formulation under this selection of uncertainty set under the distributional reward setting. This highlights that our proposed MaxMin RLHF provides a socially fair and distributionally robust solution to alignment problem with diverse user groups.
2. General Utility Reinforcement Learning: Next, we highlight the connection of our proposed formulation to general utility reinforcement learning as proposed in (Zhang et al., 2020, 2021). Zhang et al. (2020) first proposed the notion of general utility reinforcement learning that goes beyond the standard setting of cumulative reward maximization to more general utilities where standard notions of Bellman’s equation, the value function, etc fail. Thus in our formulation with diverse user preferences, once we have learned the reward models $r_{\phi_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the general utility objective ${G}_{u}({\pi})$ as the minimizer over the value-functions of user groups $u$ as $G_{u_{\texttt{min}}}(\pi):=\min_{u}\langle\rho_{\pi},r_{u}\rangle-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]}$. where, $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3). With this, the general utility formulation boils down to
$\max_{\pi}\quad$ $G_{u_{\texttt{min}}}(\pi),$ (21)
Hence we can utilize the analysis developed in (Zhang et al., 2020, 2021) for the general utility RL to derive the theoretical convergence results for our proposed algorithm in this work.
5 Experimental Results
In this section, we present a comprehensive empirical evaluation of the alignment impossibilities and our proposed solutions for language models, structured into two distinct subsections: Small Scale experiments (Sec. 5.1) for initial proof of concept, and Large Scale experiments (Sec. 5.2) for broader validation. We first demonstrate the practical challenges of alignment (cf. Theorem 3.3), followed by showcasing the efficacy of our MaxMin-RLHF strategy. This approach illustrates that, with a focus on social welfare objectives, alignment across diverse human preferences is attainable.
5.1 Small Scale Experiments (with GPT-2): Sentiment and Conciseness Alignment
Figure 5: (Alignment with MaxMin RLHF). This figure shows the performance of our proposed MaxMin RLHF algorithm for the preference dataset described in Figure 2. The task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature. The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right). The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right). We note that MaxMin-RLHF aligned language model can generate highly positive sentiment sentences and satisfy the conciseness criteria. This shows alignment with both the majority and minority preferences.
Dataset. For the experiment in this section on controlled sentiment generation, we categorized the humans into two groups: majority (Group 1) and minority (Group 2). In these sub-groups, Group 1 prefers responses with positive sentiment, and Group 2 prefers brevity (conciseness) in responses. We use the IMDb dataset as a basis for our inputs (Maas et al., 2011), the goal for the optimal policy is to produce responses y that exhibit positive sentiment (catering to Group 1) while remaining concise (catering to Group 2). We generated two sets of preference pairs for a controlled evaluation for each user group. For Group 1, we utilized a pre-trained sentiment classifier to ensure $p(\text{positive}\mid\textbf{x},\textbf{y}_{1})>p(\text{positive}\mid\textbf{x},\textbf{y}_{2})$ and similarly for Group 2 we preferred shorter responses over longer ones. To illustrate the majority and minority group dynamics, we control the proportion of the user groups in the preference data (Group 1: 80% and Group 2 - 20%). For the experiments in this subsection, we use GPT-2 (Radford et al., 2019) as the base model.
Impossibility Results. To demonstrate our impossibility results as stated in Theorem 3.3, we perform the three steps of RLHF (described in (Christian, 2020; Ouyang et al., 2022b)) as prevalent currently with a single utility reward function on the combined preference dataset. For SFT, we fine-tune GPT-2 until convergence on reviews from the train split of the IMDB dataset and use this GPT-2 backbone for both the reward model and PPO training. The generations are evaluated against the ground truth rewards $r_{1}^{*}$ for positive sentiment (majority group) and $r_{2}^{*}$ for conciseness (minority group). It is evident from Figure 3 that the generated responses are significantly biased toward the majority user group’s preference who preference positive sentiment (note high sentiment score (green curve, high score is better) after alignment) while the preferences (concise responses) of the minority user group were neglected (note high conciseness score (red curve, lower score is better) after alignment), resulting in more verbose generations than desired.
Proposed MaxMin RLHF. Our proposed algorithm can efficiently align to both group preferences as shown in Figure 5 thereby generating responses that are of positive sentiment and concise and thus cater to both the majority and minority user groups mitigating the social disparity. We further collectively present the average performance of MaxMin RLHF with the single reward RLHF and baseline model in Figure 6.
Figure 6: This figure shows the average performance in terms of sentiments of the generated output and the conciseness alignment.We note that MaxMin RLHF is able to better cater to both the alignment criteria as compared to single reward RLHF as expected.
5.2 Large Scale Experments (with Tulu2-7B)
Datasets and Experimental Setup. We use the same dataset as Jang et al. (2023) and $10$k data points from GPT4-Alpaca (Peng et al., 2023) are used as the instruction dataset to generate rollouts, collect pairwise feedback data, and PPO training. We utilize GPT-4 to simulate human annotators with preference prompts described in Table 4 in Appendix F. We divide the datasets into groups of human users. Each group has $40$ users, which are split into $30$ users in training data and $10$ users in testing data. For the experiments in this subsection, we use Tulu2-7B (Ivison et al., 2023) as the base model. For each dataset, P1, P2, and P3, we mix the training user groups to build the simulation dataset. We have 60 users in training data which are mixed from two different groups with diverse preferences. The original distribution is that users are evenly distributed in two clusters. Then, we use the EM algorithm to train $|\mathcal{U}|=2$ reward models until we converge. Update $\phi_{u},u=1,\cdots,|\mathcal{U}|$ by minimizing the negative log-likelihood loss (2). Then, trained model is used to assign clusters to users in testing data.
5.2.1 Main Results
Impossibility of Single Reward Model. When the user groups are biased in-terms of the representation of the user-groups (divided into majority and minority groups based on the preference dataset), the single reward model fails to capture the preferences of minority user groups. We test on preference dataset P1A/P1B representing two user groups and adjust the ratio of the number of users from group P1A and group P1B. Table 1 summarizes the accuracy for the majority group and minority group, as well as the accuracy of the total data. Here, low evaluation accuracy is an indication that the eventual alignment with the minority user group will be poor after the PPO step since the reward model itself is not accurate.
Ratio Total Majority Minority
1:1 0.686 0.668 0.704
2:1 0.608 0.728 0.488
6:1 0.588 0.724 0.452
10:1 0.568 0.716 0.42
Table 1: This table presents the test accuracy of the single reward model training on the preference dataset and shows its failure to align with the minority. The first column denotes the user group ratio in the dataset, the second column shows the total accuracy, the third column shows the accuracy of the majority group, and the fourth column shows the accuracy of the minority group.
Reward Learning with EM (Algorithm 2). Following the procedures from Algorithm 2) in the experiment setup, we learn a mixture of reward models with EM algorithm on all three datasets until convergence, as shown in Figure 7. From the results in Figure 7, we note that after the fourth iteration, all users are clustered correctly, meaning the mixture preference model successfully converges we successfully learn diverse groups of users with diverse preferences.
MaxMin RLHF Alignment. We evaluate the performance of our MaxMin-RLHF alignment against conventional single-reward RLHF trained on preference datasets with diversity and representational bias. Our baselines include ratios of 1, 2, 6, and 10, the same setting as discussed in Table 1 for illustration of the representational bias in preference data. Following Jang et al. (2023), we use the same $50$ instances from Koala evaluation(Geng et al., 2023) and test the model’s ability to generate answers in different groups of users’ preferences. We run pairwise evaluations by GPT-4 using AlpacaFarm codebase(Dubois et al., 2023) and use the win rate to the base Tulu2-7B model as the metric. Our results in Table 2 and Table 3 show that MaxMin alignment keeps a high win-rate in comparison with conventional RLHF methods trained by PPO with a single reward model. Our MaxMin RLHF approach achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups. Table 2 illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, whereas, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
Method P3A P3B Average
MaxMin 57.78 55.56 56.67
1:1 55.85 52.62 54.24
2:1 55.56 48.89 52.23
6:1 58.06 46.67 52.37
10:1 56.00 45.00 50.50
Table 2: Pairwise win rate (%) on P3 dataset using GPT-4.
Method P1A P1B Average
MaxMin 57.50 60.00 58.75
1:1 56.00 51.97 53.99
2:1 57.78 44.00 50.89
6:1 54.81 48.00 51.41
10:1 55.11 45.08 50.10
Method P2A P2B Average
MaxMin 54.50 56.00 55.25
1:1 53.73 54.00 53.87
2:1 55.55 51.72 53.64
6:1 52.14 49.40 50.77
10:1 53.96 45.98 49.97
Table 3: Pairwise winrate (%) on P1-P2 using GPT-4. The table illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, favoring alignment with the majority. In contrast, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
6 Conclusions
In this work, we critically examine the limitations of the conventional single-reward Reinforcement Learning from Human Feedback (RLHF) framework, particularly its insufficiency in addressing the diversity of human preferences, leading to an impossibility result for alignment with diverse preferences. To achieve a socially fair alignment in diverse human preference settings, we introduce a novel MaxMin-RLHF approach, which learns a max-min policy over mixture of reward functions to achieve a more equitable model alignment. Our experiments demonstrate the effectiveness of MaxMin-RLHF in producing socially fairer outcomes, highlighting the need for more inclusive strategies in RLHF methodologies.
Appendix A Notations
We define the various notations in this table first.
Notations Description
x prompt
$\mathcal{X}$ set of prompts
y output text generated by the LLM
$\pi_{\text{ref}}$ direct supervised fine-tuning model, takes x as input and generates y as output
$(\textbf{y}_{1},\textbf{y}_{2})$ output pair generated by LLM
$h$ human
$\mathcal{D}$ dataset which has the data of the form $(\textbf{x},\textbf{y}_{1},\textbf{y}_{2})$
$\phi$ reward model parameter
$\theta$ language model parameter
$\mathcal{H}$ set of human population
Appendix B A Detailed Context of Related Works
Reinforcement Learning from Human Feedback. RL methods, such as policy gradient, applied to train language models for long-form generation Cho et al. (2018). The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017). The PPO algorithm helps to learn a model that produces responses that maximize the reward (Ouyang et al., 2022b; Bai et al., 2022a). Besides PPO, DPO (Direct Preference Optimization, Rafailov et al. (2023)) directly trains the large language model using human preferences without training the reward model. A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework. However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023). A number of theoretical studies have analyzed the efficiency and benefits for reinforcement learning using preference data Ji et al. (2023b); Zhang et al. (2023); Li et al. (2023). Chakraborty et al. (2024) proposed a bilevel reinforcement learning framework for policy alignment. Recently (Santurkar et al., 2023) created a dataset for evaluating the alignment of language models with $60$ US demographic groups over a wide range of topics and found substantial misalignment between a selanguage models and those groups. It emphasizes the criticality of considering diversity while performing alignment.
Diversity in Human Preferences. Here, we briefly review the literature highlighting the reasons for diversity in the context of LLMs. Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a). The key factors contributing to this diversity include (i) socio-demographic backgrounds, including race, ethnicity, age, and gender shape preferences. Gender differences, for example, influence sensitivity to online content, with women facing more online harassment Vogels (2021). (ii) Personal bias and context subjectivity, which affects the human preferences for controversial topics in interpreting language and divisive themes (Denton et al., 2021b; Sandri et al., 2023)). (iii) Imperfect preferences, which arises due to variations in expertise, training, or quality control leading to diverse preferences, with certain content inaccurately considered offensive by some groups (Sandri et al., 2023). (iii) Linguistic ambiguity & missing context, could lead to diversity because of words or phrases with multiple possible interpretations and without clear context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022). These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback.
Appendix C Preliminary Results
We present the following preliminary results in the form of Lemma C and Lemma C.
{lem}
The parametrized preference probability distribution
$p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}$
under the Bradley -Terry model (Bradley & Terry, 1952) is Lipschitz with respect to parameter $\phi$. This implies that
$|p_{\phi}-p_{\phi^{\prime}}|\leq L_{p}\|\phi-\phi^{\prime}\|,$ (22)
with $L_{p}=4D$ where $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$.
Proof.
Let us start from the definition of $p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ given by
$p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ $=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}=\frac{1}{1+\exp(-(r_{\phi}({\mathbf{y}_{1}},\mathbf{x})-(r_{\phi}({\mathbf{y}_{2}},\mathbf{x}))}.$ (23)
From the definition of the Bradley-Terry preference model from equation (1) with the linear parametrization of the reward function as $r_{\phi}({\mathbf{y}},\mathbf{x})=\langle\phi,\psi({\mathbf{y}},\mathbf{x})\rangle$, we can write the equality in (23) as
$p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ $=\frac{1}{1+\exp(-(\langle\phi,\psi({\mathbf{y}_{1}},\mathbf{x})\rangle-\langle\phi,\psi({\mathbf{y}_{2}},\mathbf{x})\rangle)))}$
$=\frac{1}{1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle)},$ (24)
where we define $\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x}):=\psi({\mathbf{y}_{1}},\mathbf{x})-\psi({\mathbf{y}_{2}},\mathbf{x})\rangle$ for the ease of notation. Next, differentiating both sides in (C) with respect to $\phi$, we obtain
$\nabla_{\phi}p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ $=-\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\cdot\frac{\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle)}{(1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle))^{2}}$
$=-\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\left[\frac{1}{1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle)}-\frac{1}{(1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle))^{2}}\right].$ (25)
Taking the norm on both sides and applying Cauchy-Schwartz inequality, we get
$\|\nabla_{\phi}p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})\|$ $\leq\|\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\|\left[\frac{1}{1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle)}+\frac{1}{(1+\exp(-\langle\phi,\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\rangle))^{2}}\right]$
$\leq 2\|\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\|.$ (26)
From the definition of $\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})$ and the boundedness of the feature representations, we note that $\|\psi^{\prime}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{x})\|=\|\psi({\mathbf{y}_{1}},\mathbf{x})-\psi({\mathbf{y}_{2}},\mathbf{x})\rangle\|\leq 2D$. Hence, we obtain the final bound
$\|\nabla_{\phi}p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})\|$ $\leq 4D.$ (27)
Hence proved.
{lem}
The cross-entropy loss minimization for reward learning in step 2 in the RLHF pipeline (cf. (2)) leads to implicit weightage minimization among the user groups. Specifically, the loss function minimizes the distance to distribution $p_{\phi^{*}}({\mathbf{z}})=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{u}^{*}({\mathbf{z}})$, where $\eta$ is the implicit distribution among user groups.
Proof of Lemma C.
From the equality in (3.1), we note that we can write $p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\mathbb{E}_{u}[p^{*}_{u}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})]$. With this notation, the loss function for reward learning in (3.2) can be written as
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})$ $=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\Big{[}\mathbb{E}_{u}[p^{*}_{u}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})]\log p_{\phi}(\succ)+\mathbb{E}_{u}[p^{*}_{u}(\mathbf{y}_{1}\prec\mathbf{y}_{2}\mid\mathbf{x})]\log p_{\phi}(\prec)\Big{]},$ (28)
where the equation incorporates the individual user group’s optimal $p_{u}^{*}$ (we denote the corresponding individual optimal reward parameter by $\phi_{u}^{*}$) in the likelihood objective. As a first step, let us decompose (28) as
$\mathcal{L}_{R}$ $(r_{\phi},\mathcal{D})$
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})}\Bigg{[}\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)]\log p_{\phi}(\succ)-\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})]$
$\hskip 45.5244pt+\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)]\log p_{\phi}(\prec)-\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})]$
$\hskip 48.36967pt+\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})+\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\Bigg{]},$ (29)
where, we add and subtract $\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})]$ and $\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})]$ to get the final expression. After rearranging the terms in (C), we get
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=$ $-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\Bigg{[}\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)]\Big{(}\log p_{\phi}(\succ)-\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\Big{)}$ (30)
$\hskip 71.13188pt+\sum_{u=1}^{|\mathcal{U}|}[p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)]\Big{(}\log p_{\phi}(\prec)-\log p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\Big{)}$
$\hskip 85.35826pt+\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})$
$\hskip 99.58464pt+\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\Bigg{]}$
$=-\mathbb{E}_{x,y_{1},y_{2}}\Bigg{[}\sum_{u=1}^{|\mathcal{U}|}\eta(u)\Big{(}p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\cdot\log\frac{p_{\phi}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})}{p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})}+p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\cdot\log\frac{p_{\phi}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})}{p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})}\Big{)}$
$\ \ \ +\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}|{\mathbf{x}})+\sum_{u=1}^{|\mathcal{U}|}p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\eta(u)\log p_{u}^{*}({\mathbf{y}}_{1}\prec{\mathbf{y}}_{2}|{\mathbf{x}})\Bigg{]}.$
Next, by utilizing the definition of KL-divergence and entropy to get the final expression as follows
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})$ $=\mathbb{E}_{(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2})\sim\mathcal{D}}\Bigg{[}\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})+\eta(u)\mathsf{H}(p_{u}^{*})\Bigg{]}.$ (31)
From the above objective in (31), we note that the objective is minimized for $\phi$ when $\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})=0.$ To proceed further, let us focus on the term $\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})$ from equation (31) as
$\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})$ $=\sum_{u=1}^{|\mathcal{U}|}\eta(u)\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log\frac{p_{u}^{*}({\mathbf{z}})}{p_{\phi}({\mathbf{z}})}$
$=\sum_{u=1}^{|\mathcal{U}|}\eta(u)\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{u}^{*}({\mathbf{z}})-\sum_{u=1}^{|\mathcal{U}|}\eta(u)\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})$
$=-\sum_{u=1}^{|\mathcal{U}|}\eta(u)H(p_{u}^{*})-\sum_{\mathbf{z}}\log p_{\phi}({\mathbf{z}})\underbrace{\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{u}^{*}({\mathbf{z}})}_{=p^{*}({\mathbf{z}})}.$ (32)
From the definition of KL d in (3.1), it holds that
$\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})$ $=-\sum_{u=1}^{|\mathcal{U}|}\eta(u)H(p_{u}^{*})-\sum_{\mathbf{z}}p^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}}).$ (33)
Next, by adding and subtracting the term $\sum_{\mathbf{z}}p^{*}({\mathbf{z}})\log p^{*}({\mathbf{z}})$ in the right hand side of (33), we get
$\sum_{u=1}^{|\mathcal{U}|}\eta(u)\mathsf{KL}(p_{u}^{*}||p_{\phi})$ $=-\sum_{u=1}^{|\mathcal{U}|}\eta(u)H(p_{u}^{*})-\sum_{\mathbf{z}}p^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})+\sum_{\mathbf{z}}p^{*}({\mathbf{z}})\log p^{*}({\mathbf{z}})-\sum_{\mathbf{z}}p^{*}({\mathbf{z}})\log p^{*}({\mathbf{z}})$ (34)
$=-H(p^{*})-\sum_{u=1}^{|\mathcal{U}|}\eta(u)H(p_{u}^{*})+\mathsf{KL}(p^{*}||p_{\phi}).$
Now, replacing this expression in the original implicit minimization objective in (31), we note that the minimization will be achieved when $p_{\phi^{*}}({\mathbf{z}})=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{u}^{*}({\mathbf{z}})$ for all $z$. Hence, the reward learning objective is implicitly learning a weighted combination, which would lead to a significant gap in individual utilities, as discussed in the subsequent section. ∎
Appendix D Proof of Lemma 3.2
Proof.
Let us reconsider the reward learning loss $\mathcal{L}_{R}(r_{\phi},\mathcal{D})$ whose empirical version is minimized to obtain parameter $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter
$\phi^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{\phi^{*}}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}$
. As discussed in Sec. 3.2, due to human user groups, a user group specific $\phi^{*}_{u}$ will also exist. Our goal is to characterize the gap between $\hat{\phi}_{\texttt{MLE}}$ and $\phi^{*}_{u}$ defined as
$\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*},$ (35)
where the optimal $\phi_{u}^{*}$ for the user group $u$ is given by
$\phi_{u}^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}.$ (36)
Let us consider the idealistic setting of infinite data under which we know that MLE would converge to optimal $\phi^{*}$ (Zhu et al., 2023). Hence, to proceed further, let us add subtract $\phi^{*}$ in the right-hand side of (35), we get
$\Delta_{u}^{r}=\underbrace{\hat{\phi}_{\texttt{MLE}}-\phi^{*}}_{=0}+{\phi^{*}-\phi_{u}^{*}}.$ (37)
To derive the lower bound on the reward suboptimality $\Delta_{u}^{r}$, we begin with the definition of the total variation distance as
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\frac{1}{2}\sum_{\mathbf{z}}|p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi^{*}}({\mathbf{z}})|.$ (38)
From the Lipschitzness of the preference probability as derived in Lemme C, we can write
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\leq 2D\|\phi_{u}^{*}-\phi^{*}\|.$ (39)
From the lower bound in (38) and the expression in (37), we obtain
$\|\Delta_{u}^{r}\|\geq\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}}).$ (40)
Next, we expand on the term TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$. From the statement of Lemma C, we note that $p_{\phi^{*}}({\mathbf{z}})=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{\phi_{u}^{*}}({\mathbf{z}})$, hence we can write
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi^{*}}({\mathbf{z}}))$
$=\sum_{\mathbf{z}}\Big{(}p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)},$ (41)
where we use $k$ to denote a user group for ease of notation. Since $\sum_{k=1}^{|\mathcal{U}|}\eta(k)=1$, we can write
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}\Big{(}\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)}.$ (42)
After interchanging the order of summation, we get
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\Big{(}\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi_{k}^{*}}({\mathbf{z}}))\Big{)}$
$=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (43)
Using the equality in (D) into the right hand side of (40), we obtain
$\|\phi^{*}-\phi_{u}^{*}\|\geq\frac{1}{2D}\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (44)
From the Definition 2, we will get the final result. Hence proved.
Appendix E Proof of Theorem 3.3
Proof.
We can define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $u$ as
[…]
We note that in this specific RLHF setting under the KL-based regularization, the objective $-F_{{\phi}}(\pi)$ satisfies strong convexity w.r.t $\pi$ with strong convexity parameter $\mu=1$, hence it holds that
[…]
Now utilizing that $\log(\pi({\mathbf{y}}/x))$ is Lipschitz continuous with parameter $L_{\pi}=\frac{1}{c}$, under the condition that there exists some $c>0$ such that $\pi(y|x)\geq c$ for all $x,y$, we get
[…]
From the results in (Rafailov et al., 2023), we can derive an analytical mapping from reward functions to optimal policies for the KL-constrained reward maximization objective as denied in (10) as :
$\pi_{r}(y\mid x)=\frac{1}{Z(x)}\pi_{\text{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r(y,x)\right)$ (48)
where $\pi_{r}$ is the optimal policy under the reward $r$ and $Z(x)$ is the partition function given as $Z(x)=\sum_{y}\pi_{\text{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r(y,x)\right)$. Note that such an equivalence is specific to the RLHF problem under the Bradley Terry preference model as shown in (Rafailov et al., 2023). Next, replacing equation (48) in the equation (47), we get
[…]
$=\frac{1}{2L_{\pi}\beta^{2}}\|\Psi,\langle\phi^{*}-\phi^{*}_{u},\rangle\|^{2}.$
As stated in (C), under the linearly parametrized reward function, we have $r_{\phi}(\mathbf{y},\mathbf{x})=\langle\phi,\psi({\mathbf{y}},\mathbf{x})\rangle$, where the parameter $\phi\in\mathbb{R}^{d}$ and similarly $\psi({\mathbf{y}},\mathbf{x})\in\mathbb{R}^{d}$. Let $(x,y)\in\mathbb{R}^{n}$ and we denote the feature matrix $\Psi\in\mathbb{R}^{n\times d}$ as $\begin{bmatrix}\Psi^{T}=\Psi(y_{1},x_{1})&\Psi(y_{2},x_{2})&\cdots&\Psi(y_{n},x_{n})\end{bmatrix}$, replacing in (49), we get the final expression. Next, expanding the norm on the right hand side, we obtain
[…]
Next we lower-bound the matrix norm of $\Psi^{T}\Psi\in\mathbb{R}^{d\times d}$ with the minimum eigen value $\lambda_{\psi}$ as
[…]
where we obtain the lower bound in terms of the reward suboptimality. From the statement of Lemma 3.2, we can lower bound the right hand side in (51) as follows
[…]
Hence proved.
Appendix F Additional Details of the Experiments
In this section, we provide additional details of the experiments in Section 5.
Table 4: Dataset Summary
User Group Preference Prompt
P1A Generate/Choose a response that can be easily understood by an elementary school student.
P1B Generate/Choose a response that only a PhD Student in that specific field could understand.
P2A Generate/Choose a response that is concise and to the point, without being verbose.
P2B Generate/Choose a response that is very informative, without missing any background information.
P3A Generate/Choose a response that is friendly, witty, funny, and humorous, like a close friend.
P3B Generate/Choose a response (that answers) in an unfriendly manner.
Figure 7: Results on Dataset P1A/P1B.

### v2

@@
-MaxMin-RLHF: Towards Equitable Alignment of Large Language Models with Diverse Human Preferences
+MaxMin-RLHF: Alignment with Diverse Human Preferences
 Abstract
 Reinforcement Learning from Human Feedback (RLHF) aligns language models to human preferences by employing a singular reward model derived from preference data.
-However, such an approach overlooks the rich diversity of human preferences inherent in data collected from multiple users.
+However, the single reward model overlooks the rich diversity of human preferences inherent in data collected from multiple users.
 In this work, we first derive an impossibility result of alignment with single reward RLHF, thereby highlighting its insufficiency in representing diverse human preferences.
-To provide an equitable solution to the problem, we learn a mixture of preference distributions via an expectation-maximization algorithm and propose a MaxMin alignment objective for policy learning inspired by the Egalitarian principle in social choice theory to better represent diverse human preferences.
-We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution.
-We present comprehensive experimental results on small-scale (GPT-2) and large-scale language models (with Tulu2-7B) and show the efficacy of the proposed approach in the presence of diversity among human preferences.
-Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach.
+Next, we propose to learn a mixture of reward models via an expectation-maximization algorithm and solve a MaxMin alignment objective inspired by the Egalitarian principle in social choice theory to better honor diverse human preferences.
+We present comprehensive experimental results on small-scale (GPT-2) and large-scale language (with Tulu2-7B)) and show the efficacy of the proposed approach in the presence of diversity among human preferences.
 We remark that our findings in this work are not only limited to language models but also extend to reinforcement learning in general.
+Machine Learning, ICML
+[…]
+[…]
 1 Introduction
-The alignment problem, central to developing and fine-tuning current large language models (LLMs), represents a crucial challenge in artificial intelligence, especially in ensuring these models operate in harmony with human values, preferences and social welfare (Wang et al., 2023; Christian, 2020).
+The alignment problem, central to developing and fine-tuning current large language models (LLMs), represents a crucial challenge in artificial intelligence, especially in ensuring these models operate in harmony with human values and preferences (Wang et al., 2023; Christian, 2020).
 Reinforcement learning from human feedback (RLHF) has emerged as a pivotal approach to alignment problems, specifically aligning LLM (Wang et al., 2023; Ouyang et al., 2022b; Stiennon et al., 2022a; Ouyang et al., 2022a).
+RLHF starts with pre-training a generative (LLM) model and subsequently fine-tuning it through supervised learning on a high-quality dataset on various downstream tasks.
 RLHF operates in three steps (a) supervised fine-tuning, (2) reward learning, and (3) RL fine-tuning.
@@
 (2023a, b), “the notion of ‘one truth’ in crowdsourcing responses is a myth” and we need to account for the diversity in opinions and preferences.
-Despite the criticality, most of the latest RLHF approaches ignore the consideration of the diversity in human preference feedback by aligning the language model with a single reward function (Wang et al., 2023; Christian, 2020; Stiennon et al., 2022a; Ouyang et al., 2022a).
-This assumption of a single ground truth reward is restrictive and can potentially subdue the preferences or opinions of minority groups, leading to societal biases (Figure 1).
-To mitigate this issue, some of the recent research proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners (Bakker et al., 2022; Jang et al., 2023).
+Despite the criticality, most of the latest RLHF approaches ignore the consideration of the diversity in human preference feedback by aligning the language model with a single reward (Wang et al., 2023; Christian, 2020; Stiennon et al., 2022a; Ouyang et al., 2022a).
+The assumption of a single ground truth reward is restrictive and can potentially subdue the preferences or opinions of minority groups, leading to societal biases (Figure 1).
+To mitigate this issue, some of the recent research proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners (Bakker et al., 2022).
 On the other hand, (Ovadya, 2023) adopts a consensus-based method for aggregating human representations by emphasizing specific principles (Bai et al., 2022b; Kovač et al., 2023), which might result in the under-representation of marginalized groups (Ramé et al., 2023).
 Another line of research focuses on the aspect of designing multi-policy strategies by fine-tuning personalized language models towards individual rewards (Jang et al., 2023; Ramé et al., 2023; Ji et al., 2023a).
-Some recent literature has brought attention to the challenge of aligning single utility RLHF with diverse preferences (Bakker et al., 2022; Ramé et al., 2023; Jang et al., 2023).
+As mentioned above, the recent literature has brought attention to the challenge of aligning single utility RLHF with diverse preferences.
 However, a thorough understanding of how the diversity within human sub-populations influences the overall alignment objective remains elusive.
-Consequently, this prompts us to pose the following question:
-Is the current single reward RLHF pipeline sufficient to align with diverse human preferences?
-In this work, we present negative results for the above question by demonstrating the impossibility of alignment using single reward RLHF (Theorem 3.3).
+Consequently, this prompts us to pose the following question: Is a single reward RLHF pipeline sufficient to align with diverse human preferences?
+In this work, we present negative results for the above question in this work by demonstrating the impossibility of alignment using single reward RLHF (Theorem 3.3).
 We introduce a notion of diversity between human subpopulations due to the differences in preference distributions and establish lower bounds on the alignment performance of single reward RLHF.
 However, this impossibility result naturally raises another important question:
-What strategies can we design to efficiently align with diverse human preferences?
-In response to this question, we draw inspiration from the Egalitarian rule (Sen, 2017) to provide an equitable solution to the problem and aim to maximize the social utility objective for alignment.
-We propose learning a mixture of preference distributions via expectation-maximization algorithm in absence of human sub-population group information.
-Under the mixture distributions, we introduce the notion of MaxMin RLHF, for policy learning inspired by the Egalitarian principle in social choice theory to better represent human preferences under diversity.
-We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution.
-We summarize our contributions as follows.
+What strategies can we design (or what methods can we adopt) to align with diverse human preferences?
+In response to this question, we draw inspiration from the Egalitarian rule (Sen, 2017) and aim to maximize the social utility objective for alignment. We summarize our contributions as follows.
 Figure 1: This figure highlights the drawbacks of a single reward-based current state-of-the-art alignment framework called Reinforcement Learning from Human Feedback (RLHF) (Christian, 2020).
@@
 (1) An impossibility result of alignment with single reward-based RLHF.
-We first introduce the notation of diversity as a divergence between the human sub-population preference distributions (Definition 2) and derive lower bounds on the reward model sub-optimality (Lemma 3.2) in terms of the diversity.
+We first introduce the notation of diversity (Definition 2) and then derive lower bounds on the reward model suboptimality (Lemma 3.2) in terms of diversity in human sub-population preference distributions.
 Finally, we establish a lower bound (Theorem 3.3) on the alignment gap due to the diversity in the human preference feedback.
-True to our knowledge, ours is the first work to formally characterize the sub-optimality gap with conventional RLHF methods under diverse preferences.
+True to our knowledge, our work is the first to report such a result in the RLHF literature.
 (2) Max-Min RLHF alignment with diverse user preferences.
-To provide an equitable solution to the problem, we learn a mixture of preference distributions using the Expectation-Maximization (EM) algorithm (Algorithm 2) and propose a novel MaxMin-RLHF algorithm as an egalitarian strategy to align language models with social utility objectives under diverse preferences.
-(Algorithm 1).
+We propose to learn a mixture of preference distributions through the application of multiple reward functions using the Expectation-Maximization (EM) algorithm (Algorithm 2).
+Upon obtaining multiple reward functions specific to different human sub-populations, we introduce the MaxMin-RLHF algorithm as a strategy to align language models with social utility objectives (Algorithm 1).
 (3) A comprehensive empirical study.
-We present a detailed empirical analysis of our proposed approach on both small (GPT-2) and large language models (Tulu-7B).
+We present a detailed empirical analysis of our proposed concepts on two language models: GPT-2 and Tulu-7B.
 Initially, we provide empirical evidence highlighting the impossibilities of alignment with single reward RLHF, followed by demonstrating the feasibility and effectiveness of MaxMin-RLHF in achieving social utility objectives.
-Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach.
-1.1 Related Works
-The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017).
-Besides PPO, DPO (Direct Preference Optimization, Rafailov et al.
-(2023)) directly trains the large language model using human preferences without training the reward model.
-A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework.
-However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023).
-Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a).
-The key factors contributing to this diversity include (i) socio-demographic backgrounds (Vogels, 2021), (ii) personal bias and context subjectivity (Denton et al., 2021b; Sandri et al., 2023)), (iii) Imperfect preferences, (Sandri et al., 2023), and (iv) Linguistic ambiguity & missing context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022).
-These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback.
-Appendix B provides a detailed context of related works.
+Our approach outperforms existing methodologies, showcasing significant performance improvements.
 2 Preliminaries
@@
 (2)), which fails to account for the inherent diversity among the human sub-populations (see Figure 2).
-As discussed in Section 1.1, one of the key reasons for the diverse human preferences is the varied socio-demographic and socio-cultural backgrounds of human sub-populations (Aroyo et al., 2023b, a).
+One of the key reasons (discussed in Appendix B in ) for the diverse human preferences is the varied socio-demographic and socio-cultural backgrounds of human sub-populations (Aroyo et al., 2023b, a).
 For example, population groups with diverse demographic markers such as race, ethnicity, age groups, genders, etc., have highly varied preferences as highlighted in (Aroyo et al., 2023b, a; Denton et al., 2021a).
@@
 [Diversity in Human Preferences] Consider a human population $\mathcal{H}$, composed of $|\mathcal{U}|$ sub-population groups where $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$, and a sub-population-specific preference $p^{*}_{u}$ as defined in (5), we define the diversity of sub-population group $\mathcal{H}_{i}$ with respect to other group $\mathcal{H}_{j}$ as
-$\texttt{Diverity}\ (i,j):=\text{TV}(p_{i}^{*},p^{*}_{j}),$ (6)
-where TV denotes the total variation distance between two preference distributions. By utilizing the definition of sub-population groups in $\mathcal{U}$, we can express the preference in (4) as
+$\texttt{Diversity}\ (i,j):=\text{TV}(p_{i}^{*},p^{*}_{j}),$ (6)
+where TV denotes the total variation distance between two preference distributions.
+By utilizing the definition of sub-population groups in $\mathcal{U}$, we can express the preference in (4) as
 $p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=$ $\sum_{u=1}^{|\mathcal{U}|}\bigg{[}\sum_{h\in\mathcal{H}_{u}}\mathbb{I}_{h}(\mathbf{z})\cdot q(h|u)\bigg{]}\cdot\eta(u)$
@@
 1) and $\phi^{*}_{u}$ models the human sub-population group $\mathcal{H}_{u}\in\mathcal{U}$ specific $p_{u}^{*}$, it holds that
-$\underbrace{\|\phi^{*}-\phi_{u}^{*}\|}_{\textbf{Reward mismatch}}\geq\frac{1}{2D}\cdot{\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\texttt{Diverity}\ (u,k)},$
-where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, and diversity as defined in Definition 2.
+$\underbrace{\|\phi^{*}-\phi_{u}^{*}\|}_{\textbf{Reward mismatch}}\geq\frac{\epsilon(1-\eta(u))}{4D},$ (10)
+where $\epsilon:=\texttt{Diverity}(u,j)-\max_{k\neq u}\texttt{Diverity}(k,j)>0$, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, and diversity as defined in Definition 2.
 Proof Sketch.
 Here we describe the proof sketch of Lemma 3.2 with a detailed proof provided in Appendix D.
-We begin with the definition of sub-optimality in the learned reward for a subpopulation group $u$ as $\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*}$ where $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter $\phi^{*}$ which minimizes the loss function in (2).
-However, we know in the limit of infinite data, $\hat{\phi}_{\texttt{MLE}}$ converges to $\phi^{*}$ and hence we focus on the sub-optimality gap due to diversity as $\|\phi_{u}^{*}-\phi^{*}\|$.
-Using the Lipschitzness of the preference probability distribution under the Bradley-Teryy preference model (derived in Lemma C in Appendix) we lower-bound the sub-optimality gap by $\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ and finally expanding upon the definition of $p^{*}$ as shown in (3.1), we get the final result.
+We begin with the definition of sub-optimality in the learned reward for a subpopulation group $u$ as $\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*}$ where $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter $\phi^{*}$.
+However, we know in the limit of infinite data under appropriate regulatory conditions, $\hat{\phi}_{\texttt{MLE}}$ converges to $\phi^{*}$, and hence we focus on the sub-optimality gap due to diversity as $\|\phi_{u}^{*}-\phi^{*}\|$.
+Using the Lipschitzness of the preference probability distribution under the Bradley-Tery preference model (derived in Lemma C in Appendix) we lower-bound the sub-optimality gap and finally expanding upon the definition of $p^{*}$ as shown in (3.1), we get the final result.
 Remark.
 Lemma 3.2 indicates that the current RLHF-based reward learning paradigm (Christian, 2020; Bai et al., 2022a; Rafailov et al., 2023) will suffer sub-optimality due to diversity amongst the humans, which is highly likely in practice (Aroyo et al., 2023b).
-Lemma 1 implies that the degree to which the learned reward parameter diverges from optimality for a given subgroup is influenced by two key factors: the distinctiveness of that subgroup’s preferences compared to all the other subgroups, and the relative weight assigned to the subgroup in the overall preference model.
-3.3 An Impossibility Result of Alignment
-To mathematically characterize the impossibility of aligning the language model with diverse sub-population groups, let us reconsider the RL fine-tuning optimization problem, which is given by (step 3 in RLHF in Section 2)
-$\max_{{\pi}}F_{{\phi}}(\pi),$ (10)
+Lemma 3.2 implies that the degree to which the learned reward parameter diverges from optimality for a given subgroup is influenced by two key factors: the distinctiveness of that subgroup’s preferences compared to all the other subgroups, and the relative weight assigned to the subgroup in the overall preference model.
+3.3 An Impossibility Results of Alignment
+To mathematically characterize the impossibility of aligning the language model with diverse sub-population groups, let us reconsider the RL fine-tuning optimization problem, which is given by (step 3 in RLHF)
+$\max_{{\pi}}F_{r_{\phi}}(\pi),$ (11)
 Figure 3: (Empirical Evidence of Impossibility).
-This figure provides empirical evidence of the impossibility of alignment (cf.
-Theorem 3.3) in single reward RLHF on preference dataset presented in Figure 2.
-Here, the task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature.
-The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right).
-The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right).
-We note that the aligned language model (top row right) can generate highly positive sentiment sentences but completely ignores the requirement of conciseness (generates longer sentences).
-This is happening because the humans who prefer conciseness are in minority as described in Figure 2.
-where we define $F_{{\phi}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}\Big{[}\mathbb{E}_{\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})\bigr{]}-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot~{}|~{}\mathbf{x})||\pi_{\text{ref}}(\cdot~{}|~{}\mathbf{x})\bigr{]}\Big{]}$.
-Let us define $\pi_{\text{RLHF}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}}}(\pi)$ where $\pi_{\text{RLHF}}$ is the optimal aligned policy with single reward RLHF.
-On the other hand, we define a human sub-population specific optimal policy as $\pi^{*}_{\text{u}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}_{u}}}(\pi)$, where $\pi^{*}_{\text{u}}$ is the optimal aligned policy with individual subpopulation group $\mathcal{H}_{u}$.
-We define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $\mathcal{H}_{u}$ by
+This figure validates our theoretical results in Theorem 3.3 and provides empirical evidence of the impossibility of alignment in single reward RLHF on preference dataset presented in Figure 2.
+Here, the task is to align the LLM to generate positive sentiment responses which are concise.
+We note that the aligned language model can generate highly positive sentiment sentences but completely ignores the requirement of conciseness.
+This is happening because the humans who prefer conciseness are in the minority as compared to humans who prefer a positive sentiment score as described in Figure 2.
+where we define $F_{r_{\phi}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}\Big{[}\mathbb{E}_{\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})\bigr{]}-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot~{}|~{}\mathbf{x})||\pi_{\text{ref}}(\cdot~{}|~{}\mathbf{x})\bigr{]}\Big{]}$.
+Let us define $\pi^{*}_{\text{RLHF}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{r_{\phi^{*}}}(\pi)$ where $\pi^{*}_{\text{RLHF}}$ is the optimal aligned policy with single reward RLHF.
+On the other hand, we define a human sub-population specific optimal policy as $\pi^{*}_{\text{u}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{r_{\phi^{*}_{u}}}(\pi)$, where $\pi^{*}_{\text{u}}$ is the optimal aligned policy with individual subpopulation group $\mathcal{H}_{u}$.
+We define the alignment gap of RLHF model $\pi^{*}_{\text{RLHF}}$ to a specific user group $\mathcal{H}_{u}$ by
 […]
 We note that the alignment gap defined in (12) measures the discrepancy between the reward returns by the single reward RLHF model $\pi_{\text{RLHF}}$ and the optimal model $\pi_{u}^{*}$ tailored for $\mathcal{H}_{u}$ subpopulation evaluated under true reward function $r_{u}^{*}$.
-Next, we present our impossibility result in Theorem 3.3.
-{thm}[An Impossibility Result] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf.
-1), $\phi^{*}_{u}$ denotes the human sub-population group $\mathcal{H}_{u}\in\mathcal{U}$ specific reward function to model $p_{u}^{*}$, and alignment gap is as defined in (11).
+Next, we present our impossibility result in Theorem 3.3 {thm}[An Impossibility Result] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf.
+1), $\phi^{*}_{u}$ denotes the human sub-population-group $\mathcal{H}_{u}\in\mathcal{U}$ specific reward function to model $p_{u}^{*}$, and alignment gap is as defined in (12).
 Then, it holds that
 […]
-where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, $\lambda_{\psi}$ denotes the minimum eigenvalue of the feature matrix, $\beta$ is the regularization parameter of RLHF framework, and diversity as defined in Definition 2.
+where $\epsilon:=\texttt{Diversity}(u,j)-\max_{k\neq u}\texttt{Diversity}(k,j)>0$, $\eta$ denotes the representation for the human sub-population group $u$, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, $\lambda_{\psi}$ denotes the minimum eigenvalue of the feature matrix, $\beta$ is the regularization parameter of RLHF framework and diversity as defined in Definition 2.
 A detailed proof of Theorem 3.3 is provided in Appendix E.
@@
 Remark.
-Theorem 3.3 shows that high subpopulation diversity inevitably leads to a greater alignment gap for the model $\pi_{\text{RLHF}}$ aligned with single reward-based RLHF framework.
+Theorem 3.3 shows that high subpopulation diversity inevitably leads to a greater alignment gap.
+Here, $\epsilon$ depends on the diversity among user groups, highlighting that when the diversity between a specific user group $u$ and others is significantly higher compared to inter-group diversity, it suggests that $u$ is a minority.
+Consequently, aligning to this particular user-group with single-reward RLHF becomes particularly challenging.
+Moreover, as the representation of the user-group reduces i.e $\eta(u)\to 0$, the alignment gap further increases making it harder to align to this user-group.
 In summary, if a subgroup exhibits distinctive preferences or constitutes a minority with a smaller representation, the resulting model from single reward RLHF setting cannot accurately reflect the sub-population’s specific preferences.
-We provide empirical evidence of impossibility of alignment in Figure 6.
-4 Proposed Approach: Alignment with Diverse Preferences
-From the statement of Theorem 3.3, we note that it is impossible to align the language model with diverse human preferences with a single reward model based RLHF.
-Now the natural question arises is what can we do in practice to deal with the diverse human preferences alignment problem?
-This is exactly what we answer in this section.
-We start by proposing to learn a mixture of preference distributions as follows.
+We provide empirical evidence of the impossibility of alignment in Figure 5.
+4 MaxMin-RLHF: One Possibility
+From the statement of Theorem 3.3, it is clear that it is not possible to align diverse human preferences with a single reward RLHF.
+We start by noting that even if we can bypass the sub-optimality in reward learning (cf.
+Lemma 3.2) by learning multiple reward functions $\hat{\phi}_{u}$ for all $\mathcal{H}_{u}$, it doesn’t resolve the eventually aim of language model alignment.
+This is because our goal is to develop a single model $\pi^{*}$ that honors diverse user preferences without demonstrating bias towards specific groups such as minorities.
+To achieve that, we take motivation from the Egalitarian rule in social choice theory (Sen, 2017), which states that society should focus on maximizing the minimum utility of all individuals.
+Hence, we write our proposed alignment objective which maximizes the social utility as
+$\pi^{*}_{\mathcal{F}}\in\arg\max_{\pi}\min_{u\in\mathcal{U}}F_{r_{\phi_{u}^{*}}}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (14)
+where, $F_{r_{\phi_{u}^{*}}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\big{[}{r_{\phi_{u}^{*}}}(\mathbf{y},\mathbf{x})\big{]}$ (cf.
+(3)) represents the alignment objective for the $u^{\texttt{th}}$ sub-population or group among set of humans.
 Algorithm 1 MaxMin RLHF
@@
 8:     Update policy $\pi$ towards maximizing the objective:
-9:     $\pi_{t+1}\leftarrow\text{PPO-update}(F_{{\phi_{u}^{*}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]})$
+9:     $\pi_{i+1}\leftarrow\text{PPO-update}(F_{r_{\phi_{u}^{*}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]})$
 10:  end for
 11:  Output: Policy $\pi_{T}$ aligned with socially fair preference dataset
-4.1 Learning a Mixture Distribution for Diverse Preferences
-To deal with the diversity in preferences, we first propose learning a mixture distribution for preferences rather than a single preference distribution, which is prevalent in the current RLHF pipelines (Christian, 2020; Ouyang et al., 2022b).
-The intuition behind learning a mixture of preference models is rooted in the principles of statistical learning theory (Lindsay, 1995; Seidel, 2011) which states that a combination of multiple distributions can approximate a wide range of complex, real-world distributions more effectively than a single model.
-This concept is akin to the property for mixture models, which suggests that a mixture model can approximate any number of diverse user sub-populations to an arbitrary level of accuracy, given a sufficient number of components and appropriate parameterization (Goodfellow et al., 2016).
-Thus, the preference distribution as shown in (3.1) can be expressed as
-$p({\mathbf{z}}^{\prime})=\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (13)
-where the event $\mathbf{{\mathbf{z}}}^{\prime}:=(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})$, $y_{w},y_{l}$ are the chosen and rejected response by the human sub-population-group $\mathcal{H}_{u}$.
-With the definition in (13), the log-likelihood maximization objective to learn the sub-population preference distributions $\{p_{\phi_{u}}\}_{u=1}^{|\mathcal{U}|}$ parametrized by $\Phi=[\phi_{1},\phi_{2},\cdots\phi_{|\mathcal{U}|}]$ (under the Bradley Terry preference model in (1)) and the $|\mathcal{U}|$ groups/clusters can be written as
-$\mathcal{L}(\Phi)$ $=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (14)
-$=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}\frac{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)}{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)+\exp\left(r_{\phi_{u}}(\mathbf{y}_{l},\mathbf{x})\right)}\cdot\eta(u).$
-We learn the parameter $\Phi^{*}$ by maximizing the above log-likelihood objective.
-It is important to emphasize that mixture models do not require prior knowledge of user group classifications, i.e., which user belongs to which group, and can learn the distribution of preferences in an unsupervised manner.
-This capability underscores the generality and flexibility of mixture distributions in capturing the diversity of user preferences without explicit user group information.
-Remark.
-If we have access to individual human sub-population information, we can directly learn the reward parameter via maximizing likelihood per user group.
+MaxMin RLHF.
+If we have access to individual human sub-population rewards, we can go directly to solve the optimization problem in (14) with the algorithm summarized in Algorithm 1.
 But often, in practice, they are hardly available.
@@
 We summarize the EM algorithm for reward learning in Algorithm 2.
-Figure 4: (a) This figure shows a Robotic navigation scenario where the objective of the agent is to distribute goods among the two groups denoted by green and orange boxes.
-In the above figure, (b) shows the trajectory when only green user group preferences (they want the agent to visit them) are considered to decide the agent path.
-(c) shows the trajectory when only green user group preferences are considered to decide the vehicle path.
-(d) shows the result for our proposed formulation (cf.
-Definition 4.2), where goal is to maximize the social utility, which ensures a robust solution to satisfy all user group preferences.
-4.2 MaxMin RLHF
-We begin by noting that while we circumvent the sub-optimality in reward learning (cf.
-Lemma 3.2) sub-population specific reward models ${\phi}^{*}_{u}$ for all subsets in $\mathcal{U}$ using Algorithm 2, this does not address the ultimate goal of language model alignment.
-Since the eventual objective is to develop a single policy $\pi^{*}$ that accommodates diverse user preferences while avoiding bias towards specific groups such as minorities.
-To achieve that, we take motivation from the Egalitarian rule in social choice theory (Sen, 2017), which states that society should focus on maximizing the minimum utility of all individuals.
-Hence, we write our proposed alignment objective, which maximizes the social utility as follows {defi}(Alignment with Diverse Human Preferences) From the mixture of preference distributions in (14), we get $\{r_{\phi_{u}^{*}}\}_{u=1}^{|\mathcal{U}|}$ where $\phi_{u}^{*}$ denotes reward model parameter for each human subpopulation in $\mathcal{U}$.
-We define the alignment objective in such diverse settings as
-$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\Big{(}\min_{u}\mathbb{E}\big{[}{r_{\phi_{u}^{*}}}(\mathbf{y},\mathbf{x})\big{]}\Big{)}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (15)
-where the expectation is with respect to $\{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})\}$.
-Proof of Concept: An illustration of MaxMin RLHF.
-To illustrate the social utility of the Max-Min objective within the context of RLHF, in Figure 4, we present a simplified robotic scenario.
-In this scenario, the robotics agent’s goal is to distribute goods among two user groups $u_{1},u_{2}$ located at green and orange boxes, respectively.
-Each user group would prefer that the agent visit them and distribute goods.
-Hence, we define the reward for each user group $u_{i}$ is defined as follows: the agent receives a reward $r_{u}(s,a)=C_{u}$ if it reaches the goal state for the user $u$ and $r_{u}(s,a)=0$, otherwise.
-We demonstrate that in this basic scenario, our proposed max-min algorithm ensures that the agent serves all the user groups irrespective of the difference in utility between the user groups $\Delta_{C}=|C_{1}-C_{2}|$.
-Mathematically, from the Definition 4.2, we can write
-$\pi^{*}_{\text{MaxMin}}\in\mathop{\mathrm{argmax}}_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (16)
-[…]
-$\pi^{*}_{\text{MaxMin}}\in$ $\mathop{\mathrm{argmax}}_{\pi}\min_{u}\sum_{s,a}r_{\phi_{u}}(s,a)\cdot\lambda^{\pi}(s,a)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
-$\in$ $\mathop{\mathrm{argmax}}_{\pi}{\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (17)
-where $\lambda^{\pi}$ denote the occupancy measure corresponding to policy $\pi$.
-Let us focus on the minimization objective $\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})$, where $(s_{u},a_{u})$ signifies the goal state and action for a user $u$, and $\mathcal{C}_{u}$ is the corresponding reward.
-When the policy $\pi$ does not visit the goal for user $u$, the value of $\lambda^{\pi}(s_{u},a_{u})$ is zero, making the minimum always zero.
-Therefore, the maximum objective aims to learn a policy that maximizes the reward with respect to the users $u$ until $\lambda_{u}>0$ for all users.
-Even when the difference $\Delta_{c}=|\mathcal{C}_{1}-\mathcal{C}_{2}|$ is significantly large, highlighting a disparity between majority and minority user groups, the max-min approach effectively learns a policy that respects the needs of both users, demonstrating its social utility as shown in Figures 4.
-MaxMin RLHF Algorithm.
-Next, we introduce the algorithm to solve the MaxMin RLHF proposed in Definition 4.2 in Algorithm 1.
-After learning the reward in Step 1 for individual user sub-population $r_{u}$ using Algorithm 2, we next compute the minimum utility user group as $u_{\text{min}}\leftarrow\arg\min_{u}F_{u}(\pi)$.
-Once the minimum utility user subpopulation group is determined, we fine-tune the language model policy for that group by taking gradient steps at each iteration $t$ as
-$\pi_{{t+1}}=\pi_{t}+\eta\nabla_{\pi}(F_{u_{\text{min}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]}),$ (18)
-where $F_{u_{\text{min}}}(\pi_{t}):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi_{t}(\cdot~{}|~{}\mathbf{x})}\big{[}{r_{u_{\text{min}}}}(\mathbf{y},\mathbf{x})\big{]}$ and $\eta$ denotes the step size.
-Note that for the implementation in practice, we leverage the PPO algorithm (Schulman et al., 2017) as commonly used in existing RLHF pipelines (Stiennon et al., 2022b; Christian, 2020; Ouyang et al., 2022b).
-Algorithm 2 Learning Mixture Distribution with EM Algorithm
+Algorithm 2 Learning Rewards with EM Algorithm
 1:  Input: Preference data $\mathcal{D}$, $|\mathcal{U}|$ clusters of users among all humans in $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$, pretrained $\{r_{\phi_{u}}\}_{u=1}^{|\mathcal{U}|}$, loss function $\operatorname{loss}$, convergence criteria
@@
 4:        E-step (hard cluster assignment): assign $h$ to the $u$-th cluster s.t.
-$\mathcal{H}_{u}:=\ arg\max_{u\in{1,\cdots,|\mathcal{U}|}}\prod_{(\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},h)\in\mathcal{D}}w(\phi_{u},\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}})$
+$u=\ arg\max_{u\in{1,\cdots,|\mathcal{U}|}}\prod_{(\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},h)\in\mathcal{D}}w(\phi_{u},\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}})$
 where $w(\cdot)=\frac{\exp\left(r_{\phi_{u}}(\mathbf{y_{1}},\mathbf{x})\right)}{\exp\left(r_{\phi_{u}}(\mathbf{y_{1}},\mathbf{x})\right)+\exp\left(r_{\phi_{u}}(\mathbf{y_{2}},\mathbf{x})\right)}$
@@
 7:  end while
-4.3 Connection with DRO and General Utility RL
-In this subsection, we highlight the generality of our proposed formulation beyond social utility design to broader formulations of general utility reinforcement learning (Zhang et al., 2020) and distributionally robust optimization (Duchi et al., 2018).
-We discuss then in details follows.
-1.
-Distributionally Robust Optimization: We begin by noting that once we have learned the mixture of reward models $r_{{\phi}_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the problem of learning a robust policy as a distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Duchi et al., 2018).
-In DRO, the objective is to maximize the worst-case expected utility over an uncertainty set of distributions.
-In our scenario, this boils down to constructing an uncertainty set over the distribution of rewards $r\in\mathcal{Q}$.
-Now, expanding the RL fine-tuning objective in the equation
-$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\min_{r\in\mathcal{Q}}\mathbb{E}_{r(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (19)
-where $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3).
-The uncertainty set $\mathcal{Q}$ is typically obtained by constructing a distribution across the reward distributions $r$, however, this may lead to an overtly pessimistic solution.
-Hence, recent literature (Sagawa et al., 2020) has shown leveraging the prior knowledge of user-group information, can mitigate this by constructing the uncertainty set $\mathcal{Q^{\prime}}$ in terms of these groups $\mathcal{H}_{u}\in\mathcal{U}$, also referred to as the group distributionally robust optimization (Group-DRO).
-In Group DRO, the uncertainty set over reward distribution $\mathcal{Q}$ is constructed by a mixture of $u$ groups as $\mathcal{Q^{\prime}}=\sum_{u}\delta_{u}r_{u},\delta_{u}\in\Delta_{u}$, where $\Delta_{u}$ represents the $u-1$ dimensional probability simplex of preference distributions.
-However, since the inner problem is a linear program and the optimum is attained at the vertex, the resulting group DRO problem boils down to solving
-$\pi^{*}_{\text{MaxMin}}$ $\in\arg\max_{\pi}\min_{u}\mathbb{E}_{r_{u}(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (20)
-$\in\arg\max_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
-where $r_{u}(y,x)$ represents the reward distribution of the user sub-population group $u$.
-It is important to note that the group-DRO formulation boils down to the Max-Min RLHF formulation under this selection of uncertainty set under the distributional reward setting.
-This highlights that our proposed MaxMin RLHF provides a socially fair and distributionally robust solution to alignment problem with diverse user groups.
-2.
-General Utility Reinforcement Learning: Next, we highlight the connection of our proposed formulation to general utility reinforcement learning as proposed in (Zhang et al., 2020, 2021).
-Zhang et al.
-(2020) first proposed the notion of general utility reinforcement learning that goes beyond the standard setting of cumulative reward maximization to more general utilities where standard notions of Bellman’s equation, the value function, etc fail.
-Thus in our formulation with diverse user preferences, once we have learned the reward models $r_{\phi_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the general utility objective ${G}_{u}({\pi})$ as the minimizer over the value-functions of user groups $u$ as $G_{u_{\texttt{min}}}(\pi):=\min_{u}\langle\rho_{\pi},r_{u}\rangle-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]}$.
-where, $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3).
-With this, the general utility formulation boils down to
-$\max_{\pi}\quad$ $G_{u_{\texttt{min}}}(\pi),$ (21)
-Hence we can utilize the analysis developed in (Zhang et al., 2020, 2021) for the general utility RL to derive the theoretical convergence results for our proposed algorithm in this work.
 5 Experimental Results
@@
 The task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature.
-The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right).
-The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right).
 We note that MaxMin-RLHF aligned language model can generate highly positive sentiment sentences and satisfy the conciseness criteria.
@@
 Impossibility of Single Reward Model.
-When the user groups are biased in-terms of the representation of the user-groups (divided into majority and minority groups based on the preference dataset), the single reward model fails to capture the preferences of minority user groups.
+When the user groups are biased (divided into majority and minority groups based on the preference dataset), the single reward model fails to capture the preferences of minority user groups.
 We test on preference dataset P1A/P1B representing two user groups and adjust the ratio of the number of users from group P1A and group P1B.
-Table 1 summarizes the accuracy for the majority group and minority group, as well as the accuracy of the total data.
-Here, low evaluation accuracy is an indication that the eventual alignment with the minority user group will be poor after the PPO step since the reward model itself is not accurate.
+Table 1 summarizes the accuracy for the majority group and minority group, as well as the accuracy on the total data.
+Here, low accuracy means that the alignment with the minority user group will be poor after the PPO step since the reward model itself is not accurate.
 Ratio Total Majority Minority
@@
 Reward Learning with EM (Algorithm 2).
-Following the procedures from Algorithm 2) in the experiment setup, we learn a mixture of reward models with EM algorithm on all three datasets until convergence, as shown in Figure 7.
+Following the procedures in the experiment setup, we get similar and good results on all three datasets, as shown in Figure 6.
 From the results in Figure 6, we note that after the fourth iteration, all users are clustered correctly, meaning the mixture preference model successfully converges we successfully learn diverse groups of users with diverse preferences.
-MaxMin RLHF Alignment.
-We evaluate the performance of our MaxMin-RLHF alignment against conventional single-reward RLHF trained on preference datasets with diversity and representational bias.
-Our baselines include ratios of 1, 2, 6, and 10, the same setting as discussed in Table 1 for illustration of the representational bias in preference data.
-Following Jang et al.
-(2023), we use the same $50$ instances from Koala evaluation(Geng et al., 2023) and test the model’s ability to generate answers in different groups of users’ preferences.
-We run pairwise evaluations by GPT-4 using AlpacaFarm codebase(Dubois et al., 2023) and use the win rate to the base Tulu2-7B model as the metric.
-Our results in Table 2 and Table 3 show that MaxMin alignment keeps a high win-rate in comparison with conventional RLHF methods trained by PPO with a single reward model.
-Our MaxMin RLHF approach achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups.
-Table 2 illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, whereas, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
+MaxMin RLHF Alignment. We further test the performance of our MaxMin-RLHF alignment method and compare it with the single reward
 Method P3A P3B Average
@@
 Table 2: Pairwise win rate (%) on P3 dataset using GPT-4.
-Method P1A P1B Average
-MaxMin 57.50 60.00 58.75
-1:1 56.00 51.97 53.99
-2:1 57.78 44.00 50.89
-6:1 54.81 48.00 51.41
-10:1 55.11 45.08 50.10
-Method P2A P2B Average
-MaxMin 54.50 56.00 55.25
-1:1 53.73 54.00 53.87
-2:1 55.55 51.72 53.64
-6:1 52.14 49.40 50.77
-10:1 53.96 45.98 49.97
+RLHF models trained on biased datasets.
+Our baselines include ratios of 1, 2, 6, and 10, the same setting as discussed for Table 1.
+Following Jang et al.
+(2023), we use the same $50$ instances from Koala evaluation(Geng et al., 2023) and test the model’s ability to generate answers in different groups of users’ preferences.
+We run pairwise evaluations by GPT-4 using AlpacaFarm codebase(Dubois et al., 2023) and use the win rate to the base Tulu2-7B model as the metric.
+Our results in Table 2 and Table 3 show that MaxMin alignment keeps a high win rate while the models trained by PPO with a single reward model on biased datasets will have a relatively poor performance on the minority data representing minority user groups.
+Method P1A P1B
+MaxMin 57.50 60.00
+1:1 56.00 51.97
+2:1 57.78 44.00
+6:1 54.81 48.00
+10:1 55.11 45.08
+Method P2A P2B
+MaxMin 54.50 56.00
+1:1 53.73 54.00
+2:1 55.55 51.72
+6:1 52.14 49.40
+10:1 53.96 45.98
 Table 3: Pairwise winrate (%) on P1-P2 using GPT-4.
-The table illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, favoring alignment with the majority.
-In contrast, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
 6 Conclusions
-In this work, we critically examine the limitations of the conventional single-reward Reinforcement Learning from Human Feedback (RLHF) framework, particularly its insufficiency in addressing the diversity of human preferences, leading to an impossibility result for alignment with diverse preferences.
-To achieve a socially fair alignment in diverse human preference settings, we introduce a novel MaxMin-RLHF approach, which learns a max-min policy over mixture of reward functions to achieve a more equitable model alignment.
+In this work, we critically examine the limitations of the single-reward RLHF framework, particularly its insufficiency in addressing the diversity of human preferences, leading to an impossibility result for alignment with diverse preferences.
+To achieve a socially fair alignment in diverse human preference settings, we introduce a novel approach called MaxMin-RLHF, which learns a max-min policy over a distribution of reward functions to achieve a more equitable model alignment.
 Our experiments demonstrate the effectiveness of MaxMin-RLHF in producing socially fairer outcomes, highlighting the need for more inclusive strategies in RLHF methodologies.
+Impact Statement
+The primary objective of this work is to highlight the limitations of the existing alignment techniques in representing diverse opinions and preferences.
+Our research is one of the first to formally highlight the above limitation with mathematical and empirical demonstrations.
+Finally, our research demonstrates a first step towards equitable alignment with diverse preferences.
+The findings of our work would encourage and foster further research in the domain of alignment under diversity, ensuring the current AI models is not biased towards specific minority groups.
+Acknowledgements
+Chakraborty and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
+Manocha and Bedi are supported by Army Cooperative Agreement W911NF2120076.
+Mengdi Wang acknowledges the support by NSF IIS-2107304, NSF CPS-2312093, ONR 1006977 and Genmab.
+We also thank Rui Yang and Han Zhao for pointing out a bug in the proof of Lemma 1 in the previous version.
 Appendix A Notations
@@
 Reinforcement Learning from Human Feedback.
-RL methods, such as policy gradient, applied to train language models for long-form generation Cho et al.
-(2018).
+RL methods, such as policy gradient, applied to train language models for long-form generation (Cho et al., 2018).
 The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017).
@@
 However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023).
-A number of theoretical studies have analyzed the efficiency and benefits for reinforcement learning using preference data Ji et al.
-(2023b); Zhang et al.
-(2023); Li et al.
-(2023).
-Chakraborty et al.
-(2024) proposed a bilevel reinforcement learning framework for policy alignment.
+A number of theoretical studies have analyzed the efficiency and benefits for reinforcement learning using preference data (Ji et al., 2023b; Zhang et al., 2023; Li et al., 2023).
+(Chakraborty et al., 2024) proposed a bilevel reinforcement learning framework for policy alignment.
 Recently (Santurkar et al., 2023) created a dataset for evaluating the alignment of language models with $60$ US demographic groups over a wide range of topics and found substantial misalignment between a selanguage models and those groups.
@@
 The key factors contributing to this diversity include (i) socio-demographic backgrounds, including race, ethnicity, age, and gender shape preferences.
-Gender differences, for example, influence sensitivity to online content, with women facing more online harassment Vogels (2021).
+Gender differences, for example, influence sensitivity to online content, with women facing more online harassment (Vogels, 2021).
 (ii) Personal bias and context subjectivity, which affects the human preferences for controversial topics in interpreting language and divisive themes (Denton et al., 2021b; Sandri et al., 2023)).
@@
 {lem}
-The parametrized preference probability distribution
-$p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}$
-under the Bradley -Terry model (Bradley & Terry, 1952) is Lipschitz with respect to parameter $\phi$. This implies that
-$|p_{\phi}-p_{\phi^{\prime}}|\leq L_{p}\|\phi-\phi^{\prime}\|,$ (22)
-with $L_{p}=4D$ where $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$.
+The parametrized preference probability distribution $p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}$ under the Bradley -Terry model (Bradley & Terry, 1952) is Lipschitz with respect to parameter $\phi$.
+This implies that
+$|p_{\phi}(\mathbf{z})-p_{\phi^{\prime}}(\mathbf{z})|\leq L_{p}\|\phi-\phi^{\prime}\|,$ (15)
+where $\mathbf{z}:=(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$, $L_{p}=4D$, and $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$.
 Proof.
@@
 Let us reconsider the reward learning loss $\mathcal{L}_{R}(r_{\phi},\mathcal{D})$ whose empirical version is minimized to obtain parameter $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter
-$\phi^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{\phi^{*}}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}$
-.
+$\phi^{*}:=\arg\min_{\phi}-\mathbb{E}\bigg{[}\sum_{\mathbf{z}}p_{\phi^{*}}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\bigg{]}.$
 As discussed in Sec.
-3.2, due to human user groups, a user group specific $\phi^{*}_{u}$ will also exist.
+3.2, due to the presence of diverse human user groups, a $\phi^{*}_{u}$ which, is user group specific, will also exist.
 Our goal is to characterize the gap between $\hat{\phi}_{\texttt{MLE}}$ and $\phi^{*}_{u}$ defined as
@@
 where the optimal $\phi_{u}^{*}$ for the user group $u$ is given by
-$\phi_{u}^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}.$ (36)
+$\phi_{u}^{*}:=\arg\min_{\phi}-\mathbb{E}\bigg{[}\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\bigg{]}.$ (29)
 Let us consider the idealistic setting of infinite data under which we know that MLE would converge to optimal $\phi^{*}$ (Zhu et al., 2023).
@@
 From the Lipschitzness of the preference probability as derived in Lemme C, we can write
-TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\leq 2D\|\phi_{u}^{*}-\phi^{*}\|.$ (39)
+TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\leq 4D\|\phi_{u}^{*}-\phi^{*}\|,$ (32)
+where the multiplication of $2$ comes from the fact that there are two terms in the summation in the right side of (31) (cf.
+Sec.
+3.2).
 From the lower bound in (31) and the expression in (30), we obtain
-$\|\Delta_{u}^{r}\|\geq\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}}).$ (40)
-Next, we expand on the term TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$.
-From the statement of Lemma C, we note that $p_{\phi^{*}}({\mathbf{z}})=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{\phi_{u}^{*}}({\mathbf{z}})$, hence we can write
-TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi^{*}}({\mathbf{z}}))$
-$=\sum_{\mathbf{z}}\Big{(}p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)},$ (41)
-where we use $k$ to denote a user group for ease of notation. Since $\sum_{k=1}^{|\mathcal{U}|}\eta(k)=1$, we can write
-TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}\Big{(}\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)}.$ (42)
-After interchanging the order of summation, we get
-TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\Big{(}\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi_{k}^{*}}({\mathbf{z}}))\Big{)}$
-$=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (43)
-Using the equality in (D) into the right hand side of (40), we obtain
-$\|\phi^{*}-\phi_{u}^{*}\|\geq\frac{1}{2D}\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (44)
-From the Definition 2, we will get the final result. Hence proved.
+$\|\Delta_{u}^{r}\|\geq\frac{1}{4D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}}).$ (33)
+Next, to obtain a lower bound on the term TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$, we begin with the definition of the total variation distance $TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})$ as
+$TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})$ $=\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
+$=\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi^{*}}(z)+p_{\phi^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
+$\leq\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi^{*}}(z)|+\frac{1}{2}\sum_{z}|p_{\phi^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
+$=TV(p_{\phi_{u}^{*}},p_{\phi^{*}})+TV(p_{\phi_{j}^{*}},p_{\phi^{*}}).$ (34)
+In (D), the first equality holds from the definition of TV norm distance, the second equality holds because we add subtract $p_{\phi^{*}}(z)$ inside the norm, we used triangle inequality for the third inequality, and then again utilize the definition of TV distance to write the final equality.
+After rearranging the terms in (D), we can write
+$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})\geq TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-TV(p_{\phi_{j}^{*}},p_{\phi^{*}}).$ (35)
+From the definition of $p_{\phi^{*}}(z)=\sum_{k=1}^{|\mathcal{U}|}n(k)p_{\phi_{k}^{*}}(z)$ and from the property of TV distance and Jensen’s inequality, we know that $TV(p_{\phi_{j}^{*}},p_{\phi^{*}})\leq\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$.
+This would imply that $-TV(p_{\phi_{j}^{*}},p_{\phi^{*}})\geq-\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$, which we utilize in the right hand side of (35) to obtain
+$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\geq TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$
+$=(1-\eta(u))TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}}),$ (36)
+which demonstrates the final expression of our lower bound on the term $TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$, which holds in general.
+Considering the second term in the right-hand side of (D), we can write
+$\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})\leq TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})\sum_{k\neq u}\eta(k),$ (37)
+where $TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}}):=\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})$ for a given $j$.
+From the definition of weights $\eta$, we know that $\sum_{k\neq u}\eta(k)=(1-\eta(u))$, hence we can write
+$\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})\leq(1-\eta(u))TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}}).$ (38)
+Utilizing the upper bound of (38) into the right hand side of (D), we obtain
+$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\geq(1-\eta(u))\Big{[}TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})\Big{]}.$ (39)
+In the above expression, we note from the right-hand side that the lower bound in (39) holds for all $u$ and $j$.
+This implies that the right-hand side in (39) will be either positive or negative for $(u,j)$ pairs.
+But an interesting point to note here is that if the right-hand side is lower bounded away from zero, even for one $(u,j)$ pair, that is an impossibility result, for the corresponding $u$ which is the minority user.
+To proceed, as a first step, we find the most diverse user-group ($i,j\in\mathcal{U}$) defined as $(i^{*},j^{*})$ be the pair of users with the maximum total variation distance, given by
+$(i^{*},j^{*})=\arg\max_{i,j\in\mathcal{U}}TV(p_{\phi_{i}},p_{\phi_{j}}).$ (40)
+We next define the minority user group $u^{*}$ as the group with the largest total variation distance from the remaining user groups $k\in\mathcal{U},k\neq j^{*}$, where we quantify this by computing the maximum distance to any other user-groups in the set $\mathcal{U}$.
+Assuming there exists an
+$u^{*}=\arg\max\bigg{\{}\max_{k\neq j}\text{TV}(p^{*}_{i^{*}},p^{*}_{k}),\max_{k\neq i}\text{TV}(p^{*}_{j^{*}},p^{*}_{k})\bigg{\}}.$ (41)
+Thus in this process, we select that user-group which is at maximum distance from the set $\mathcal{U}$ and denote as the minority user-group $u^{*}$.
+Assuming uniqueness in the minority group definition i.e we assume that there exists an unique user-group which is at a maximum distance to the group $j$, we have that $TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})>TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})$ for $j=j^{*}$.
+Using the equality on the right hand side of (33), we obtain
+$\|\phi^{*}-\phi_{u}^{*}\|\geq\frac{\epsilon(1-\eta(u))}{4D},$ (42)
+where $\epsilon:=TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})>0$ (for $j=j^{*}$). Hence proved.
 Appendix E Proof of Theorem 3.3
 Proof.
-We can define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $u$ as
+We can define the alignment gap of RLHF model $\pi^{*}_{\text{RLHF}}$ to a specific user group $u$ as
 […]
-We note that in this specific RLHF setting under the KL-based regularization, the objective $-F_{{\phi}}(\pi)$ satisfies strong convexity w.r.t $\pi$ with strong convexity parameter $\mu=1$, hence it holds that
+We note that in this specific RLHF setting under the KL-based regularization, the objective $-F_{r_{\phi}}(\pi)$ satisfies strong convexity w.r.t $\pi$ with strong convexity parameter $\mu=1$, hence it holds that
 […]
@@
 […]
-$=\frac{1}{2L_{\pi}\beta^{2}}\|\Psi,\langle\phi^{*}-\phi^{*}_{u},\rangle\|^{2}.$
+$=\frac{1}{2L_{\pi}\beta^{2}}\|\langle\Psi,\phi^{*}-\phi^{*}_{u}\rangle\|^{2}.$
 As stated in (C), under the linearly parametrized reward function, we have $r_{\phi}(\mathbf{y},\mathbf{x})=\langle\phi,\psi({\mathbf{y}},\mathbf{x})\rangle$, where the parameter $\phi\in\mathbb{R}^{d}$ and similarly $\psi({\mathbf{y}},\mathbf{x})\in\mathbb{R}^{d}$.
@@
 […]
-Hence proved.
+where $\epsilon:=TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})>0$. Hence proved.
 Appendix F Additional Details of the Experiments
@@
 Figure 6: Results on Dataset P1A/P1B.
+Appendix G Additional Experiments in Robotics Navigation Tasks
+In this section, we show that the proposed ideas are well extendable to reinforcement learning in general.
+We show the performance of MaxMin alignment with single reward RLHF on simple gridworld navigation in Figure 7.
+Figure 7: (a) This figure shows a GridWorld navigation scenario where say a government supported vehicle which needs to distribute goods among the two groups denoted by green and orange boxes.
+In the above figure, (b) shows the trajectory when only green user preferences are considered to decide the vehicle path.
+(c) shows the trajectory when only green user preferences are considered to decide the vehicle path.
+(d) shows the result for our proposed formulation, where our goal is to maximize the social utility, which makes sure to develop a robust solution to satisfy all user preferences.